A Bibliography of Publications about the Mach Operating System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 1.35

Title word cross-reference

1 [BRS^{+}85, BCF^{+}91, BM91, EHP94, Mit91, Uhl92, Wah90]. 1-MK [MYS^{+}93]. 11th [IEE94a]. 12th [Jam92]. 13th [IEE94b]. 14th [ACM93a]. 18th [Ano95]. 1986 [ACM86]. 1992 [Jam92].

2 [KLN91, PAO93]. 2-4 [Bit92]. ’20 [Ano93a]. 20th [Ano93a]. 21st [IEE94f]. 22nd [Ciz94]. 29-March [IEE88a]. 2nd [GHR89].

3.0 [BDMVL93a, GJ91, JCS^{+}91, LHFL93, MRGB91, RS95, WKF^{+}92]. 3.0-overview [JCS^{+}91]. 3.0/UX [RS95]. 3/DE [THKS95]. 32nd [IEE93c].

4-7 [BS91]. 4.3BSD [BL90c, BL90b]. 4th [SBC^{+}94].

5th [USE95a].

6000 [Jaf93].

’86 [ACM86]. ’89 [IEE89a, IEE89b, Voe89]. 8th [TP94].

9 [MMR91]. 90-06 [Ver90]. 90-VAPP [Bur90]. ’90s [Wil88]. ’92 [ACM92a, Bd92b, Vor92]. ’93 [SBC^{+}94]. ’94 [IEE94a, NDB94, TP94]. ’95 [IEE95a]. 9th [IEE95b, MMR91].

AAECC [MMR91]. AAECC-9 [MMR91]. abstraction [MST93]. ACCENT [Ras87]. Access [BTMD89, BJ94, Roy93]. ACM [Ciz94]. Activations [BDMVL93b, BD92a, BDMVL93a]. Adding [BD92a, BDMVL93b, BDMVL93a]. address [CLFL94, Che93, Ros94]. Addressed [CMS90, IKWS92]. Advanced [Ano89a, HS94a, TS90]. Aerospace [IEE88b]. AIX [Ras91]. AL [IEE88c, IEE90b]. algebra [MMR91]. algebraic [MMR91]. algorithm [BAA94, Saa92]. Algorithms [Ano89b, MMR91]. allocation [NUMS94]. alternative [Pat93]. Alto [IEE91]. American [Boa90]. analysis [Car93, DW95, Jaf93]. anatomy [YTC94]. Annual [Ano93a, Ano93d, Ciz94, IEE89a, IEE92b, IEE94f, Shr89]. Antilles [HS94a]. Antonio [IEE93c]. API [NUMS94]. APL [CJ91, cJmC91]. Apple [Ano92d]. Application [Ano89b, RMGB91, RSS93, RS95, EKM^{+}99, GJ92, MRGB91, Pha91, TN95]. Application-transparent [RS95]. Applications [ACM86, Ano93b, BRS^{+}85, Boa90, Ciz94, Gir94, Gol90, IEE88b, IEE92b, IEE95c, IEE95d, Mai93, BM95, CB89, CB90, Dan93, GJ94, MST94, Tok95]. applications-RT-Mach [Tok95]. applicative [GW90a]. Applied [MMR91]. Approach [BL90c, BRG^{+}89, BL90b, Tev87a]. April [ACM89, Ano92c, Gir94, IEE94f, IEE95b, USE92a, USE93a, USE93b]. Arcade [Rob94]. Architectural [ACM89, ACM92b, LHC93]. Architecture [Ano93a, BST95, CMS90, GSR93, IEE94f, Jam92, Seb91b, Tev87a, BS91, BGJ^{+}91, BGJ^{+}92, MYS^{+}93, Seb91a, VBD^{+}92, Wil88]. Architecture-independent [Tev87a]. Architectures [Ano93c, USE92a, USE93d]. Architektur [Jam92]. artificial [YT91]. Ashville [ACM93a]. ASPLOS [ACM89, ACM92b]. ASPLOS-III [ACM89]. ASPLOS-V [ACM92b]. assist [WLT93]. assistants [LMR93]. Association [USE91c].

Assurance [BL89, IEE89a]. Assuring [FM93]. Asynchronous [Jen94]. Atlanta [Ano91b, USE91b]. Audio [SBC^{+}94]. August [USE88, USE92c, USE93c, Ver90]. Australia [Ano89a]. Australian [Ano89a]. automated [Vor92]. automatically [CJ91]. Automation [IEE94c]. Autumn [Ano90b, Ano92a]. Avalon [CLNW90]. Avalon/Common [CLNW90].

B3 [E^{+}91]. BACH [GAR^{+}93]. Bad [Ver90]. balancing [Jal92, Sha91]. Baltimore [Ano95]. Bandwidth [Mai93]. Barbara [IEE95b]. Barcelona [Ano93b, NDB94]. Based [Bab90b, Chexx, Mor96, And90, Bab89, Bab90a, BBP92, BHM^{+}93, Cha94, CJMT93, DLR^{+}92, GAR^{+}93, GADV91, LBLM90, LMR93, Min95, Nan91, PRK95, SC93, THKS95, TK94, Yep92, YT91]. basis [Dri92, Tev87b]. battle [TBG^{+}87]. BBN [WGR93]. Beach [USE92b]. Belgium [Bd92b]. Bell [May88]. Bell-La [May88]. belongs [Wel91]. Benchmark [FKL91, KF93]. benchmarks [FKJ^{+}92]. benefit [NDB94]. Berkeley [Bas91]. Berlin [EHP94]. between [vRBC^{+}92]. bit [Dra91]. Bologna [TP94]. Boston [ACM89, ACM92b, IEE94d, Mai93]. Boulder [ACM93c]. boundary [Pet93]. breakpoints [Yep92]. Bruges [Bd92b]. buffer [BRG^{+}89, Ros89]. buffers [BKW94, Koo93]. build [Dea93]. Building [Chexx, CFH^{+}93a, TN95]. Burlington [USE90]. Butterfly [WGR93].

C [CFK^{+}91, CKS93, GJ92, JM92, MUI95, TI94a, USE92c]. C* [Cha94]. CA [ACM91, ACM92a, ACM93b, Ano93a, Ano93c, Ano94, GHR89, IEE89b, IEE89c, IEE89d, IEE91, IEE92a, IEE93b, IEE93d, IEE94b, IEE95a, IEE95b, USE92b]. Cache [CMS90]. cached [VBD^{+}92]. caches [IKWS92, WB92]. caching [BKW94, Roy93]. California [USE91a, USE93d, USE94, IEE88a]. Calls [Pat93]. Cambridge [USE93c]. Camelot [Spe87, SPB88]. Cannes [Ano92c]. capacity [MST93, MST94]. Caracas [Gir94]. Case [MLB^{+}97]. Casper [VBD^{+}92]. catalyst [Ras89]. Cathedral [IEE88a]. Cenju [THKS95]. Cenju-3 [THKS95]. Cenju-3/DE [THKS95]. Center [Ano95]. Centre [Ano89a]. Cern [Ver90, Ver90]. Challenge [Ciz94]. checkpointing [CN92, RS95]. Chicago [IEE92b, IEE94f, IEE95c]. chip [NUMS94]. Chorus [BCC^{+}91, Imaxx]. City [USE95a]. Clara [ACM93b]. class [CJMT93]. Client [BST95, GJ92]. client-server [GJ92]. cluster [Cha94]. clusters [RP94]. CMU [Mac91, Mac92, McD89, Tob93]. co [BJ95, ACM93c]. co-resident [BJ95]. codes [MMR91]. collector [MFY91, Ono93]. Common [Mac91, Mac92, McD89, CLNW90]. Communication [GBB93b, Bas91, BHSC98, DBRD91, GBB93a, Joh91, Lie92, Rao91]. Communications [Ano92c]. Comparative [PM18]. Comparison [Jal92, GW90a, Rao91]. Comparisons [KF93]. COMPASS [IEE89a]. compatible [SR89]. COMPCON [IEE88a, IEE89b, IEE95a]. Competitive [BGW89]. Compilers [ACM93c, SS96]. Complexity [Ciz94]. Computational [Ano90a]. Computer [ACM93b, Ano93a, Ano93d, Ciz94, Cra90, IEE88a, IEE88b, IEE89a, IEE89b, IEE90a, IEE92a, IEE92b, IEE94f, IEE94c, Ish92, BJ94, Jen94, TMJY91, TS90]. Computers [Ano89b, SS96, JM92]. Computing [Amm90, Ano92a, Bab90b, Ciz94, Gir94, HS94a, IEE93a, IEE94d, IEE95d, Jam92, USE93c, Ver90, Bab89, Bab90a, BJ95, Bry88, CLNW90, EHP94, Far89, Nan91, NCS^{+}90, RNJ^{+}90, SED^{+}89]. concepts [Bau92, DF94]. concepts-design [Bau92].

Concurrency [Bla90b, MLB^{+}97, Bla90a]. concurrent [BO96, JM92, MFY91, MBS95, TI94a]. Conference [ACM86, ACM89, ACM92a, ACM92b, ACM93b, Amm90, Ano89a, Ano90b, Ano90c, Ano91a, Ano92a, Ano93d, Ano95, Bur90, Ciz94, EHP94, Gir94, HS94b, IEE88a, IEE88b, IEE89a, IEE89b, IEE90a, IEE92b, IEE93c, IEE93d, IEE94c, IEE94d, Ish92, MSNS91, MMH93, MS95, Shr89, TP94, USE91c, USE92c, USE94, USE95b, Vor92]. Configurable [IEE94e, BHSC98, GMSS94]. CONPAR [Bur90]. Conservative [BL90c, BL90b, MFY91, Ono93]. considerations [LHC93]. Consistency [WB92, BRG^{+}89, Ros89]. construct [NCS^{+}90, RNJ^{+}90]. constructing [BHSC98]. continuations [Dea93, DBRD91, Lie92]. continuous [And90, TN95, Tok95]. Control [IEE93c, Mit91, BJ94, EKM^{+}99, GKK94, Min95, TBG^{+}87, TK94, KONT95]. controlling [MR95]. convenience [PRK95]. Convention [Ano95]. copying [Ono93]. core [Bit92]. correcting [MMR91]. correctness [BT92]. coupled [BAA94]. Coyote [BHSC98]. creation [MBS95]. critical [BT92]. Cronus [Sou97]. Cronus/Mach [Sou97]. cryptography [BCB88]. Cube [TS90]. current [JCS^{+}91]. customizable [KLM^{+}93].

Dallas [USE91c]. Dana [IEE94b]. Data [DA92, cJmC91, SGM90]. Database [She91]. databases [Red92]. DC [Ano90a, Ano90c]. debugger [CB89, CB90]. Debugging [Hov91, PLL91, Yep92]. December [ACM93a, IEE88b, IEE88c, IEE93c, IEE93e]. decentralized [Jen94]. Decision [IEE93c]. decomposition [MB93]. default [GD91]. demultiplexing [YMBM94]. Dependable [EHP94]. description [Leh89]. Design [ACM92a, Ano93b, Ano94, CLR94, For88, KD95, NUS^{+}93, TST96, UNS^{+}94, BHMR91, BHM^{+}93, Bau92, BO96, IKWS92, LHC93, OKID92, SCSK93, Wil88]. Developing [FKJ^{+}92, BCR91a, BCR91b, OT95]. Development [MYS^{+}93, ABG^{+}86, GW90b, Tev87b]. Developments [Rag92]. Device [GSR93, Dan93]. Diego [Ano93a, Ano93c, IEE93d, USE93d]. Digest [IEE89b, IEE93a, IEE95a, IEE88a]. Digital [SBC^{+}94, And90, LMR93]. discardable [Sub91]. Distributed [ACM93c, Amm90, Ano88a, Ano91b, Ano92a, Ano92b, CKS93, GGDD92, Gir94, IEE90b, IEE93d, IEE94e, IEE94b, MLB^{+}97, Nic91, Seb91b, SPB88, USE91b, USE92b, BBP92, Bau92, BM92, BGW89, BCC^{+}91, CLR94, CNTS93, CLNW90, CCGS92, DW95, For88, FKS92, Hag94, Imaxx, JR86, KF93, NCS^{+}90, Pad95, PRK95, Pha91, Ras91, RNJ^{+}90, Rob94, Sha91, Spe87, SED^{+}89, TI94a, Tev87a, WLT93, Yep92, YT91, FM93, Seb91a]. distributed-memory [BM92]. Distribution [Mil94, MGZ93]. divisible [BAA94]. Dourdan [CJ92]. DPMI [GMR93]. driven [Wen88]. driver [Min93]. Drivers [GSR93, Dan93]. DROL [TST96]. Durham [Boa90, IEE93e]. Dutch [HS94a]. Dynamic [TK94, Jal92, SZ92, Sha91].

ECOOP [TP94]. EDCC [EHP94]. EDCC-1 [EHP94]. edge [Ano92d]. edition [McD89]. effective [BFS89]. effects [IKWS92]. Efficient [BAA94, IMP94, YMBM94, CFH^{+}93a]. Eighth [MS95]. ElipSys [DLR^{+}92]. Embedding [BCB88]. Emulation [Mal91, Dra91, JCS^{+}91]. Emulator [Pat93]. Enabling [Mai93]. endpoints [YMBM94]. Engineering [BS91, Bit92]. enhanced [JM92]. entities [FKS92]. Environment [BRS^{+}85, CP97, MUI95, TST96, Bry88, Cha94, GMR93, GW90b, GADV91, Pha91, Roy93, Sha91, SED^{+}89]. Environments [ACM93c, Ano89b, Bla91, Tev87a]. Eos [GADV91]. EPEX [Bol89]. Equus [Far89]. error [MMR91]. error-correcting [MMR91]. EUROMICRO [Ano93b]. Europe [Ano92c, NDB94]. European [EHP94, TP94]. EUUG [Ano90b]. Evaluation [HCF^{+}94, KONT95, PM18, TN91, For88, FKS92, Joh91, Wen88]. evolution [Ras87]. Evolving [FL94]. examples [Bau92, DF94, Leh89]. exception [Bla88, BGH^{+}89]. exclusion [BRE92]. Executing [GSR93]. Execution [CJ91, MUI95, Cha94]. expected [Ano88a]. Experience [Ano92a, BCR91a, BCR91b, CR92b, CR92a, CFH^{+}93b, Duc91, BCF^{+}93, BHM^{+}93, BCC^{+}91, Dan93, Imaxx]. Experiences [Ano91b, MGZ93, TN95, USE91b, USE92b, JCS^{+}91]. experiment [MHP94]. Experimental [IEE90b, Pha91, FKS92]. Exploitation [cJmC91]. Exploiting [Ano89b]. exporting [ST93]. Extensible [BST95, Tof89]. Extension [KTN93]. Extensions [NYM92, NM91, Tok95]. external [KN93, Nic91, Rob94, Sub91].

Fachtagung [Jam92]. facilities [Jal92]. facility [Bla88, BGH^{+}89, Sha91, Spe87]. failures [IMP94]. fair [TS90]. Fast [Bar91, BRE92, Lie92, SCB93, MFY91, OMOP93]. Fastest [AG95]. Fault [Bab89, Bab90a, Bab90b, Chexx, EKM^{+}99, IEE93a, ACCB93, BHMR91, Nan91, RSS93, SC93]. Fault-Tolerant [Bab90b, Chexx, IEE93a, Bab89, Bab90a, EKM^{+}99, ACCB93, Nan91]. Fe [USE93a, USE93b]. features [BBP92, Joh91, TS90]. February [IEE88a, IEE89b]. Fifth [ACM92b, MSNS91]. file [Roy93, Wel91]. files [TTG^{+}87]. filesystem [LBLM90]. fine [BHSC98, BM95]. fine-grain [BHSC98]. fine-grained [BM95]. First [Ano94, EHP94]. FL [IEE88b]. FLEX [CFH^{+}93a]. Flexible [SPB88, CFH^{+}93a, KN93, WW94]. Flight [Ano90a]. Florence [Ano92b]. FME [NDB94]. follow [Ano88a]. formal [GJ94, NDB94]. foundation [ABG^{+}86, RBF^{+}89, Wah90]. Fourth [IEE88b, IEE89a, IEE89b, IEE93b, IEE94c]. framework [And90]. France [Ano90b, Ano92c, CJ92, IEE93a]. Francisco [ACM92a, IEE88a, IEE89b, IEE95a, USE94]. FreeBSD [PM18]. FT [EKM^{+}99]. FT-RT-Mach [EKM^{+}99]. FTCS [IEE93a]. FTCS-23 [IEE93a]. FTM [MHP94]. functional [FKS92]. future [GGDD92, Tev87b, TS89].

GA [Ano91b, USE91b]. Gaithersburg [IEE89a]. garbage [MFY91]. general [BJ95, OMOP93, SED^{+}89]. Generalized [JCS^{+}91]. Generation [She91, Fur94]. generational [Ono93]. Generic [SJ95]. Germany [EHP94, Jam92, Ver90]. GI [Jam92]. GI-ITG-Fachtagung [Jam92]. GI-ITG-Meeting [Jam92]. Global [Ano92c]. goals [KD95]. Gothic [BBP92]. grain [BHSC98]. grained [BM95]. Graphics [Mal91]. Group [Ano89a, Boa90]. Grove [ACM91, GHR89, IEE89d]. Guest [Hov91]. Guide [BCF^{+}93, CFH^{+}93b, Hag94, HCF^{+}94].

Hammersmith [Ano89b]. handling [Bla88, BGH^{+}89, BM91, Koo93]. Hardware [Ano93b, WLT93, GAR^{+}93]. harmonizing [Ish92]. Hawaii [HS94b, MSNS91, MMH93, MS95, Shr89]. Herrenalb [Ver90]. Heterogeneous [CP97]. HI [HS94b, MSNS91, MMH93, MS95, Shr89]. High [GHR89, Mai93, She91, SED^{+}89, BM92, MB93]. High-Bandwidth [Mai93]. High-Performance [She91, BM92, MB93]. highly [LBLM90]. highly-parallelized [LBLM90]. Hill [IEE88a]. history [SC93]. HoME [OKID92]. Hotel [IEE88a].

Huntsville [IEE88c, IEE90b]. hybrid [Ono93].

I/O [FGB91b, FGB91a]. IBM [Jaf93, McD89]. Identification [BST95]. IEEE [IEE88a]. IFAC [Bd92b]. IFIP [Gir94]. II [WGR93, Ano91b, HS94b, IEE89d, USE91b]. III [ACM89, USE92b, USE93a, USE93b]. image [WGR93]. impact [CB93]. implement [DBRD91]. Implementation [ACM92a, Ano94, Bar91, MUI95, Mil94, PC90, Sha91, She91, THKS95, TST96, Wen88, BS91, For88, FKS92, Jal92, LHFL93, OKID92, OMOP93, TI94a]. implementations [GW90a]. Implementing [CB89, CB90, Red92, TNML93]. improvements [BCF^{+}91]. Improving [PRK95]. In-kernel [LHFL93]. Independent [USE93c, Tev87a]. indexed [WB92]. indivisible [BAA94]. industrial [Fur94, NDB94]. InfoJapan’90 [Ish92]. Information [Ano95, IEE93d, IEE95a, Ish92]. infrastructure [GKK94]. Institute [HS94a]. Integer [KLN91]. Integrated [And90, IEE94c, NKAT93]. integration [Sou97]. Integrity [IEE89a]. Intellectual [IEE89b, IEE88a]. intelligence [YT91]. intelligent [PHY^{+}92]. interactive [MR95]. interface [BTG^{+}88, KN93, MR95, Min93, Nic91, Rob94, TTG^{+}87]. Interfaces [Hov91, GJ91]. interim [Spe87]. Internals [Bit90]. International [ACM89, ACM92b, Amm90, Ano93a, Bur90, CJ92, GHR89, HS94b, IEE88a, IEE89b, IEE90a, IEE91, IEE92b, IEE93a, IEE93d, IEE94f, IEE94c, IEE94d, IEE94e, IEE95b, IEE95d, MMR91, MSNS91, MMH93, MS95, NDB94, SBC^{+}94, Shr89, Vor92]. Internet [Spe87]. interpretation [May88]. interprocess [Bas91, Joh91, Rao91]. interrupt [SCB93]. Introduction [Rob90, Bol89]. inversion [NKAT93]. investigation [WGR93]. invocations [GMSS94]. IP [And90, CJMT93, Moy93]. IPC [Bar91, KTN93, Koo93, MLB^{+}97, OMOP93]. Iris [Mal91]. ISCA [Ano93a]. ISMM [Amm90]. issues [BTM88]. Italy [Ano92b, TP94]. ITG [Jam92]. IV [Bur90].

January [Ano90c, HS94b, IEE93d, MSNS91, MMH93, MS95, Shr89, USE91c, USE94, USE95b]. Japan [IEE95d, Ish92]. Java [MKT98]. Joint [Bur90]. July [Ano89a, TP94, Vor92]. June [ACM92a, Ano89b, Ano91a, Bd92b, IEE89a, IEE93a, USE95a].

Kailua [Shr89]. Kailua-Kona [Shr89]. Kauai [MSNS91]. Kernel [ABG^{+}86, Ano93c, CRRS93, JR86, Leh89, RH91, TST96, TBG^{+}87, USE92a, USE93d, BHMR91, BTG^{+}88, Bit92, BTMD89, Car93, CLR94, Dri92, GMSS94, GD91, LHFL93, May88, Ras89, RJO^{+}89, RT90, SR89, Wel91, WKF^{+}92, Pet92]. kernelized [DA92]. Kernels [USE92a, BM95, CPW93, LHC93, PLL91, SCB93]. Kiel [Jam92]. Kona [Shr89]. KTK [GMSS94].

Lake [USE95a]. Lancaster [SBC^{+}94]. Language [ACM92a, BO96, CFK^{+}91, GW90a, JR86]. Languages [ACM86, ACM89, ACM92b, ACM93c, IEE90a, SS96, Ono93]. Large [BRS^{+}85, CR92b, CR92a, Koo93, Ros89, YMBM94]. Large-Scale [BRS^{+}85, CR92b, CR92a, Ros89]. latency [Jef94]. lecture [BS91, Bit92]. Level [GSR93, SP91b, Dea93, DW95, OT94, OT95, SP91a, TNML93]. Leverage [IEE89b, IEE88a]. leverage/COMPCON [IEE88a]. Libraries [Ano89b]. library [Dea93, Moy93]. Life [Pet93]. Linkage [Ros94]. Linux [Bro97]. Lisp [CLNW90, Mac91, Mac92, McD89]. load [Jal92, MGZ93, Sha91]. Local [Mil94]. Location [USE93c]. Location-Independent [USE93c]. lock [Car93, CPW93]. Logic [Vor92, DLR^{+}92]. lookaside [BKW94, BRG^{+}89, Ros89]. loosely [BAA94]. Lottery [WW94]. Louisiana [USE95b]. Low [Ros89, DW95]. Low-synchronization [Ros89]. LPAR [Vor92].

M [SGM90]. MA [ACM89, ACM92b, IEE94d, Mai93, USE93c]. Maarten [HS94a]. Mac [PM18]. Mach [AKST93, EKM^{+}99, KONT95, Ras89, TN91, USE91a, USE93a, USE93b, ABG^{+}86, ACCB93, BTG^{+}88, JR86, Lac91, Ras87, Tev87b, TBG^{+}87, TTG^{+}87, And90, Bab89, Bab90a, Bab90b, Bacxx, BCB88, BRS^{+}85, Bar91, BD92a, BDMVL93b, BDMVL93a, Bas91, Bau92, BB93, Bit90, BS91, Bit92, Bla88, BGH^{+}89, Bla90a, Bla90b, BGJ^{+}91, Bla91, BGJ^{+}92, Bol89, BCC^{+}91, BL90c, BL90b, BKLL93, BTM88, BL89, Bro97, BCCR91, Car93, CLR94, CNTS93, CB89, CB90, Cha94, CR92b, CR92a, CMS90, CJMT93, Chexx, Dan93, Dan94, Dra91, Dra92, Dri92, DW95, Duc91, EKM^{+}99, ES90, FM93, FKJ^{+}92, FL94, For88, FGB91b, FGB91a, GBB93a, GBB93b, Gol90, GD91, GJ91, Hov91, Imaxx, Jaf93, Jal92, Joh91, JCS^{+}91, KF93, KTN93, Koo93, Kup93, KD95, LBLM90, Leh89, LHFL93, MRGB91, Mal91, McD89]. Mach [MR95, MUI95, MGZ93, MZDG93, Mil94, MLB^{+}97, Min95, Mit91, Mit93, MKT98, Mor96, Moy93, NM91, NYM92, NKAT93, Nan91, Nic91, OMOP93, Pad95, PRK95, Pat93, Pet92, PAO93, Rao91, RBF^{+}89, RJO^{+}89, RT90, RMGB91, Ras91, Red92, Rob94, RSS93, RS95, Saa92, SR89, ST93, Seb91b, Seb91a, SP91a, SP91b, Sha91, Sou97, Spe87, SJ95, THKS95, TST96, Tev87a, TS89, TN95, Tof89, Tok95, USE90, Uhl92, WKF^{+}92, Wie92, Yep92]. Mach-1 [BRS^{+}85]. Mach-Based [Mor96, Cha94, LBLM90, PRK95]. Mach-US [SJ95]. Mach/4.3BSD [BL90c, BL90b]. Mach/EPEX [Bol89]. Mach/IBM [McD89]. Machine [Cra90, cJmC91]. Macintosh [Bro97]. making [Ano95]. managed [NUS^{+}93, UNS^{+}94]. Management [SP91b, BGW89, BM91, BFS89, DBRD91, Jef94, Joh91, Lie92, MR95, MDRK93, NKAT93, Nic91, Rob94, RSS93, SP91a, SCB93, Tev87a, Uhl92, WW94, WB92]. manager [GD91]. Managing [Ano92c, Sub91, MST93]. Manual [Ano93d, BTG^{+}88, Mac91, Mac92, McD89]. Manufacturing [IEE94c]. mapped [MDRK93, TTG^{+}87]. March [Ano91b, Ano92c, Ano93d, Ciz94, IEE88a, IEE89b, IEE90a, IEE94e, IEE95a, Jam92, USE91b, USE92b]. Marketplace [Ano92c]. Markov [Saa92]. Maryland [Ano95]. Masix [CLR94]. Matching [BM95]. Matchmaker [JR86]. May [ACM93b, Ano93a, IEE89c, IEE92a, IEE94d, IEE94a, IEE95c, SS96]. MC88200 [Mal91]. MD [IEE89a]. measure [FKJ^{+}92]. Measurement [ACM93b]. measurements [Dan93, Dan94, Leh89]. Measuring [CPW93]. mechanisms [BHMR91]. media [And90, TN95, Tok95]. meditation [BTMD89]. Meeting [Ano92c, Ciz94, Jam92]. Melbourne [Ano89a]. Memory [ACM93c, BB93, Bit90, BCCR91, SP91b, BM92, BGW89, BCF^{+}91, BJ94, BFS89, BM95, CR92b, CR92a, CB93, CRRS93, For88, GD91, IKWS92, JM92, Joh91, cJmC91, MDRK93, NUMS94, Nic91, Pad95, Red92, Rob94, RP94, Ros89, Saa92, SGM90, SP91a, TI94a, Tev87a, TTG^{+}87, WLT93]. memory-mapped [MDRK93]. message [BTMD89, Tof89]. messages [Koo93, YMBM94]. method [FKS92]. methods [Che93, GJ94, NDB94]. Mexico [USE93a, USE93b]. Micro [USE92a, BHMR91, BM95, Car93, CLR94, Dri92]. micro-kernel [BHMR91, Car93, CLR94, Dri92]. Micro-Kernels [USE92a, BM95]. Microkernel [BGJ^{+}91, BGJ^{+}92, CN92, MZDG93, Mil94, THKS95, BCF^{+}93, BO96, CNTS93, CJMT93, GMR93, KD95, LMR93, MGZ93]. Microkernel-based [THKS95, LMR93]. Microkernels [Ano93c, USE93d, vRBC^{+}92]. Microprocessing [Ano93b]. microprocessor [GAR^{+}93]. microprocessor-based [GAR^{+}93]. microprocessors [KLN91]. Microprogramming [Ano93b]. Microsoft [Ano92d]. microtasking [GW90a]. midrange [AG95]. Midwest [Ano93d]. Migrating [FL94]. Migration [MZDG93, MDP^{+}00, Pha91, WGR93]. MIKE [CNTS93]. MIMD [cJmC91]. MIPS [MYS^{+}93]. mission [BT92]. MITRE [GJ94]. MK [MYS^{+}93]. MKM [Leh89]. MMU [Mal91]. Mobile [USE93c]. Model [FL94, May88, Saa92, TS89, WKF^{+}92]. Modeling [ACM93b, AKST93]. modern [BM95]. module [Mal91]. monitor [GAR^{+}93, Tof89, Leh89]. monitoring [MR95, PHY^{+}92, TI94b]. Monterey [Ano94, USE91a]. move [Far89]. movement [DA92]. Moving [GD91]. MSD [GW90b]. Multi [BJ94, CJMT93, GJ91]. multi-class [CJMT93]. multi-server [GJ91]. Multi-view [BJ94]. multicast [vRBC^{+}92]. multicomputer [Roy93]. Multimedia [IEE94d, NM91, NYM92, Dan93, MST94, Dan94]. Multimedia/Realtime [NYM92, NM91]. multiple [CCGS92, GMR93, NUMS94, YMBM94]. multiple-API [NUMS94]. multiplication [KLN91]. Multiprocessor [Ano91b, BRS^{+}85, SZG92, USE91b, USE92b, BAA94, CPW93, CR92b, CR92a, JM92, PLL91, Ros89, SZ92, TI94a]. Multiprocessors [ACM93c, WWT89]. multithreaded [CB89, CB90, Pha91]. mutual [BRE92]. MVM [GMR93].

Napa [IEE93b]. Nashville [Ano91a]. National [Ano95]. NATO [HS94a]. NATUG [Boa90]. NATUG-2 [Boa90]. NC [ACM93a, Boa90, IEE93e]. nCUBE [MUI95]. need [KLM^{+}93]. Nemesis [Ros94]. Netherlands [Ano92a]. Network [Bar91, SBC^{+}94, Bas91, BL90a, OMOP93, Pet93, Ras87, RH91, TNML93]. networking [MB93]. Networks [CP97]. neural [BL90a]. Newport [USE92b]. Next [Fur94, She91, Cra90, Min93, TMJY91, TS90, Wil88]. Next-Generation [She91]. Nice [Ano90b]. Nineteenth [Ano93b]. North [Boa90]. NOSSDAV [SBC^{+}94]. notes [BS91, Bit92]. November [Ano90a, Ano92a, Ano94, BS91, Bit92, SBC^{+}94, USE91a]. NT [AG95]. NUMA [BFS89]. NY [Amm90, IEE94c, SS96].

O [FGB91b, FGB91a]. O.S. [Hov91]. Oakland [IEE89c, IEE92a]. Object [ACM86, CJ92, Dri92, GJ91, HCF^{+}94, IEE91, SJ95, TP94, BCF^{+}93, BBP92, BHM^{+}93, BCC^{+}91, CNTS93, CFH^{+}93b, GADV91, Hag94, Imaxx, JR86, KD95, MFY91, Min95, Min93, Ono93, TS89]. object-based [BBP92, GADV91]. Object-Orientated [ACM86]. Object-Oriented [TP94, GJ91, BCF^{+}93, BCC^{+}91, CNTS93, Hag94, Imaxx, JR86, KD95, MFY91, Min93, Ono93, TS89]. Objects [BST95, GMSS94, MDRK93]. OCCAM [Ano89a]. October [ACM86, ACM91, ACM92b, ACM93c, Amm90, Ano90b, Ano95, Boa90, EHP94, HS94a, IEE90b, IEE91, IEE93b, IEE94c, IEE94b, IEE95d, Ish92, MMR91, NDB94, USE90]. OlDiLa [Bau92]. on-chip [NUMS94].

OOPSLA [ACM86]. Open [Ano93b, Ras89, RBF^{+}89, Wah90]. OpenForum [Ano92a]. Operating [ACM89, ACM91, ACM92b, ACM93a, Ano94, BRS^{+}85, BT92, Bla90b, BCR91c, CJ92, IEE89d, IEE91, IEE93b, IEE94a, MBS95, MHP94, RP94, SBC^{+}94, THKS95, Tok95, Ano88a, BCF^{+}93, Bla90a, BGJ^{+}91, BGJ^{+}92, BJ95, BL89, BCR91a, BCR91b, CPW93, CLR94, CLFL94, CB93, CCGS92, DBRD91, Dri92, DW95, FKJ^{+}92, FKS92, Fuk93, Fur94, GGDD92, Jal92, Jef94, Joh91, KLM^{+}93, Lac91, LMR93, MST94, MRZ94, NUMS94, NM91, PLL91, Rag92, Ras87, RBF^{+}89, Ras91, Rob90, Ros94, SR89, Sha91, Sta94, SCB93, TS89, Tof89, WWT89]. operations [Min95]. OptiMach [Bacxx]. Optimal [NUMS94, Ben92]. Optimistic [Bacxx]. optimization [Koo93]. Orientated [ACM86]. Orientation [CJ92, IEE91]. Oriented [TP94, BCF^{+}93, BCC^{+}91, CNTS93, Dri92, GJ91, Hag94, Imaxx, JR86, KD95, MFY91, Min93, Ono93, TS89]. Orlando [IEE88b]. Orleans [IEE90a, MMR91, USE95b]. OS/2 [PAO93]. OS/network [Pet93]. OSDI [Ano94]. OSF [Bit92, BCF^{+}91, BM91, Mit91, MYS^{+}93, Uhl92, Wah90]. OSF/1 [Bit92, BCF^{+}91, BM91, Mit91, MYS^{+}93, Uhl92, Wah90]. OSF/1-MK [MYS^{+}93]. Other [Ano93c, USE92a, USE93d]. Overview [Seb91a, FKL91, JCS^{+}91].

PA [Ano88b, IEE94e]. Pacific [ACM91, GHR89, IEE89d]. packages [OT95]. packet [YMBM94]. Padula [May88]. Page [Dra91, SC93, Saa92]. pager [Sub91]. pages [Sub91]. paging [KN93]. Palo [IEE91]. Papers [IEE89b, IEE93a, IEE95a, IEE88a]. paradigm [CCGS92]. Parafrase [YTC94]. Parafrase-2 [YTC94]. Parallel [Amm90, Ano89b, Ano92b, Bur90, CP97, Fuk93, Gir94, IEE93d, IEE95b, Nil92, THKS95, BL90a, Bry88, BCR91c, CFK^{+}91, DF94, DLR^{+}92, MFY91, TI94b, Tev87a, Tob93]. Parallelism [Ano89b, Bla90b, Bla90a, cJmC91]. Parallelization [BL90c, BL90b]. parallelized [CJ91, LBLM90]. Parallelizing [BM91]. PARSAC [KLN91]. PARSAC-2 [KLN91]. passing [BTMD89]. PC [Cha94, CFK^{+}91]. Performance [BL90a, Dan93, Dan94, GHR89, Jaf93, Joh91, Rao91, She91, BM92, Car93, CPW93, CB93, FKJ^{+}92, For88, FKS92, IKWS92, KF93, LHFL93, MB93, SED^{+}89, TS90]. persistence [VBD^{+}92]. Persistent [HCF^{+}94, CRRS93, MDRK93]. personal [LMR93]. Personality [PAO93]. Petersburg [Vor92]. Phoenix [Ciz94]. Physical [SP91b, SP91a]. Physics [Cra90]. Pittsburgh [Ano88b, IEE94e]. platform [CNTS93]. PMAP [Mal91]. Point [IEE94b]. policy [Bla91, Min95]. Port [Koo93]. porting [CR92b, CR92a]. Portland [ACM86, USE88, USE92c]. Practice [Ano92a]. prefetching [BKW94, SC93]. Preprints [Bd92b]. present [GGDD92]. Primitives [GBB93b, GBB93a]. Principles [ACM91, ACM93a]. priority [Bla91, CJMT93, NKAT93, SCB93]. priority-based [CJMT93]. Privacy [IEE89c, IEE92a]. problem [Ben92]. Proceeding [MMH93]. Proceedings [ACM86, ACM89, Amm90, Ano88b, Ano89a, Ano90b, Ano90c, Ano91a, Ano92a, Ano92c, Ano93c, Ano93d, Ano94, Boa90, Bur90, CJ92, Ciz94, GHR89, HS94a, HS94b, IEE88c, IEE89a, IEE89c, IEE89d, IEE90b, IEE91, IEE92a, IEE92b, IEE93b, IEE93c, IEE93e, IEE93d, IEE94f, IEE94c, IEE94d, IEE94a, IEE94e, IEE94b, IEE95b, IEE95c, IEE95d, MSNS91, MS95, SBC^{+}94, Shr89, TP94, USE88, USE91a, USE91c, USE92a, USE92c, USE93a, USE93b, USE93c, USE93d, USE94, USE95b, Ver90, Ano95, EHP94, Ish92, MMR91, NDB94, USE90, Vor92]. Process [IEE89a, MDP^{+}00, BM91]. Processing [Bur90, IEE95b, SPB88, Nil92, SGM90, SED^{+}89, Tob93, WWT89]. Processor [MST93, MST94, Wen88]. Processors [Bla91]. production [YT91]. Programming [ACM86, ACM89, ACM92a, ACM92b, Bd92b, BKLL93, CP97, TP94, BO96, BCR91c, CFK^{+}91, CNTS93, CCGS92, DF94, DLR^{+}92, JM92, NCS^{+}90, RNJ^{+}90, TI94b, Vor92]. programs [Cha94, CJ91, GMR93, YTS88, Yep92]. proportional [WW94]. proportional-share [WW94]. prospective [Fur94]. Protection [Hag94, BM95, CLFL94, MRZ94]. Protocol [MB93, RH91]. protocols [TNML93]. Prototype [E^{+}91, PC90, WKF^{+}92]. Providing [Cha94, Min95]. purpose [BJ95, SED^{+}89].

QOS [TK94, KONT95]. QoS-Control [KONT95].

R [GKK94]. R-TICS [GKK94]. Raleigh [IEE93e]. reactive [PHY^{+}92]. reader [Ben92]. reader-writer [Ben92]. Real [BB93, Bd92b, Ciz94, HS94a, IEE88c, IEE93e, IEE94a, IEE95c, IEE95d, KONT95, KTN93, MKT98, ST93, SZG91, Sta94, TST96, Ano95, BJ95, Dan93, EKM^{+}99, Fur94, GKK94, LHC93, MRZ94, NCS^{+}90, OT94, OT95, RNJ^{+}90, SZG92, SZ92, TN91, TK94, AKST93, Dan94, NKAT93, TN95]. Real-Time [Bd92b, IEE88c, IEE93e, IEE94a, IEE95c, IEE95d, KONT95, KTN93, MKT98, TST96, ST93, SZG91, Sta94, BJ95, Dan93, EKM^{+}99, Fur94, GKK94, LHC93, MRZ94, NCS^{+}90, OT94, OT95, RNJ^{+}90, SZG92, SZ92, TN91, TK94, AKST93, Dan94, NKAT93, TN95]. Real-World [Ciz94]. Realtime [NYM92, DW95, Jen94, NM91]. reasoning [Vor92]. Rechensystemen [Jam92].

Recoverable [Pad95, CRRS93]. recoverable-persistent [CRRS93]. Recovery [Bacxx, Gol90, Red92]. Redirecting [Pat93]. reference [Dra91]. refined [May88]. refinement [Bau92]. Reflection [OT95]. Related [HCF^{+}94]. Reliable [CLNW90, IEE94b, vRBC^{+}92, BHMR91, BHM^{+}93, NCS^{+}90, RNJ^{+}90]. Remote [MLB^{+}97, MBS95]. replacement [Dra91, Saa92]. Report [Mit93, Spe87]. requirement [BT92]. Research [Boa90, IEE92a, Nil92]. Researches [Tob93]. reserves [MST93, MST94]. resident [BJ95, SGM90]. Resource [Mit91, MR95, WW94]. results [MHP94]. review [AG95]. RIG [Ras87]. RISC [Jaf93]. RP3 [Bry88, BCR91a, BCR91b, BCR91c, CR92b, CR92a, CJ91]. RPC [Duc91, IMP94]. RT [EKM^{+}99, EKM^{+}99, KTN93, McD89, MR95, Tok95]. RT-IPC [KTN93]. RT-Mach [EKM^{+}99, MR95]. RTOSS [IEE94a]. Run [ACM93c, MDRK93, SS96]. Run-Time [ACM93c, SS96, MDRK93]. running [GMR93, Yep92]. Runtime [TST96]. Russia [Vor92].

Safety [IEE89a]. Salt [USE95a]. San [ACM92a, Ano93a, Ano93c, IEE88a, IEE89b, IEE93c, IEE93d, IEE95a, USE93d, USE94]. Santa [ACM93b, IEE95b, USE93a, USE93b]. Scalable [SS96]. Scale [BRS^{+}85, CR92b, CR92a, Ros89, WWT89]. Scaling [Ciz94]. Scheduler [BDMVL93b, BD92a, BDMVL93a, AKST93]. Scheduling [Bla90a, Bla90b, WWT89, BAA94, Bla91, PHY^{+}92, WW94, Wen88]. schizophrenic [SCSK93]. School [Ver90]. Science [Ciz94]. Sciences [HS94b, MSNS91, MMH93, MS95, Shr89]. Seattle [IEE94a, USE92a]. Second [CJ92, IEE89d, IEE93d, IEE94e, IEE95d, NDB94, Shr89]. secure [Ben92]. securing [YTS88]. Security [Ano95, BTM88, IEE88b, IEE89a, IEE89c, IEE92a, USE88, USE95a]. SEDMS [Ano91b, USE91b, USE92b]. self [YTS88]. self-securing [YTS88]. semi [Saa92]. semi-Markov [Saa92]. September [ACM86, ACM93c, Ano88b, Ano92b, Ano93b, Ano93c, Bur90, CJ92, GHR89, IEE89d, IEE92b, Mai93, USE93d, Ver90]. Server [BST95, MKT98, ACCB93, Bas91, For88, GJ91, GJ92, ES90]. Servers [KONT95, SJ95, BHMR91, Dan93, LHFL93, RH91]. service [CJMT93, MB93]. services [BHSC98, JCS^{+}91, Nic91]. Seventh [Ano93d, HS94b]. share [WW94, Ano92c]. Shared [BCCR91, HCF^{+}94, BGW89, CR92b, CR92a, CFH^{+}93b, For88, Jef94, JM92, cJmC91, Pad95, Rob94, RP94, Ros89, TTG^{+}87, WLT93]. shared-memory [CR92b, CR92a, cJmC91, Ros89]. Sharing [CLFL94]. SICS [Nil92]. Sigmetrics [ACM93b]. signal [BM91]. SIGPLAN [ACM92a]. Silicon [Mal91]. Simple [BFS89]. simulations [BL90a]. single [CLFL94, Ros94]. single-address-space [CLFL94]. Sint [HS94a]. SISAL [GW90a]. Sixteenth [IEE92b]. Sixth [MMH93]. Sizing [TS90]. small [Koo93, WWT89]. small-scale [WWT89]. Smalltalk [OKID92]. Society [IEE88a, IEE89b, IEE92a, Ish92]. Software [Ano89b, Ano93b, BKW94, Che93, HS94b, IEE89a, IEE92b, IEE94a, Shr89, TMJY91, Uhl92, Wah90, BRG^{+}89, GW90b, NUS^{+}93, RJO^{+}89, RT90, TI94b, Tof89, UNS^{+}94, Voe89, WGR93]. software-managed [NUS^{+}93, UNS^{+}94]. solution [Ben92]. Some [BBP92]. sound [Min93]. space [CLFL94, Ros94]. Spain [Ano93b, NDB94]. Spring [Ano92c, IEE88a, IEE89b]. Sprite [Kup93]. SR [BO96]. St [Vor92]. Stardust [CP97]. status [JCS^{+}91]. Step [Bau92]. stock [KLN91]. storage [MDRK93]. StrongBox [YTS88]. structure [CB93]. Structured [BCCR91]. Study [HS94a, MLB^{+}97]. Subjects [BST95]. Suite [FKL91, KF93]. suited [BCF^{+}93]. Summer [Ano91a]. Supercomputer [WGR93]. Supercomputers [Ano88b]. superdatabase [PC90]. Superhighway [IEE95a]. Support [ACM89, ACM92b, Bla90b, BJ95, HCF^{+}94, Hov91, SBC^{+}94, BCF^{+}93, BO96, Bla90a, BCR91c, CN92, CFH^{+}93b, CRRS93, GMSS94, JR86, MST94, MR95, MDRK93, MBS95, Red92, RH91, RP94, Tok95, YTS88, Yep92]. Supporting [BCC^{+}91, BCCR91, Imaxx, BJ95, VBD^{+}92]. Switzerland [Bur90]. Symposium [ACM91, ACM93a, Ano90a, Ano91b, Ano93a, Ano93b, Ano93c, Ano94, IEE88c, IEE89c, IEE92a, IEE93a, IEE93e, IEE94f, IEE94b, IEE95b, IEE95c, MMR91, Mit93, NDB94, USE91a, USE91b, USE92b, USE93a, USE93b, USE93c, USE93d, USE95a]. synchronization [Ros89, TN91]. synthesis [WGR93]. System [Ano93b, Bla90b, Chexx, ES90, FGB91a, HCF^{+}94, HS94b, Jaf93, MSNS91, MMH93, MS95, Pat93, SBC^{+}94, Shr89, SPB88, THKS95, Ano88a, BBP92, BHM^{+}93, BCB88, BT92, BHSC98, Bla90a, BGJ^{+}91, BGJ^{+}92, BJ94, BCC^{+}91, BTM88, BL89, BCR91a, BCR91b, BCR91c, CPW93, CLR94, CLFL94, CB93, Che93, CFH^{+}93b, DLR^{+}92, Dri92, DW95, FKJ^{+}92, FGB91b, FKS92, GJ91, Hag94, Imaxx, Joh91, Lac91, LMR93, MST94, MBS95, Min95, Moy93, MHP94, NM91, PLL91, PRK95, PHY^{+}92, Rag92, Ras87, RJO^{+}89, RT90, RP94, Ros94, SR89, Sha91, SCB93, SCSK93, TS89, Tof89, Tok95, Wel91, WWT89, YT91, E^{+}91, SGM90]. Systems [ACM86, ACM89, ACM91, ACM92b, ACM93b, ACM93a, Amm90, Ano91b, Ano92b, Ano94, Ano95, GHR89, IEE88c, IEE89a, IEE89d, IEE90b, IEE91, IEE93b, IEE93e, IEE93d, IEE94d, IEE94a, IEE94e, IEE94b, IEE95d, Jam92, She91, SS96, USE91b, USE92b, BCF^{+}93, BAA94, Bau92, BJ95, CJ92, CFH^{+}93a, CCGS92, DA92, DF94, DBRD91, Fuk93, Fur94, GGDD92, GAR^{+}93, GADV91, GKK94, IMP94, Jal92, Jef94, Jen94, JR86, KLM^{+}93, MFY91, MRZ94, NUMS94, Ras89, RBF^{+}89, Ras91, Rob90, Sta94]. systems-supporting [BJ95]. systems/Unix/AIX [Ras91].

Task [MZDG93, MBS95]. Tasking

[MLB^{+}97]. Tasks [Bacxx, GSR93, BAA94]. TC2000 [WGR93]. TCP [And90, CJMT93, Moy93]. TCP/IP [And90, CJMT93, Moy93]. Technical [Ano92a, USE92c, USE95b]. Techniques [She91, BFS89]. Technologies [IEE95a, Mai93]. Technology [Ano89a, Ano90a, HS94b, IEE94c, IEE95c, Voe89, BCF^{+}93, Ish92, Rag92, TMJY91]. Telecommunications [Ano89a]. Temporal [MRZ94]. testbed [Dan93, SGM90, Dan94]. Third [ACM89, IEE88a, IEE93a]. Thirteenth [ACM91]. Thirty [IEE88a, IEE89b]. Thirty-Fourth [IEE89b]. Thirty-Third [IEE88a]. Thread [DF94, FL94, DBRD91, Lie92, OT95]. Threads [Duc91, NCS^{+}90, RNJ^{+}90, Dea93, OT94, SZG91, SZG92, SZ92, TI94b, TBG^{+}87, TK94]. Three [GJ94]. TICS [GKK94]. Time [ACM93c, Bd92b, HS94a, IEE88c, IEE93e, IEE94a, IEE95c, IEE95d, KONT95, KTN93, MKT98, SS96, TST96, BJ95, Dan93, EKM^{+}99, Fur94, GKK94, Jef94, LHC93, MRZ94, MDRK93, NCS^{+}90, OT94, OT95, RNJ^{+}90, ST93, SZG91, SZG92, SZ92, Sta94, TN91, TK94, Wen88, AKST93, Dan94, NKAT93, TN95]. time-driven [Wen88]. time-shared [Jef94]. timers [ST93]. TLB [Uhl92]. TLBs [NUS^{+}93, UNS^{+}94]. TMach [May88]. TN [Ano91a]. Tokyo [IEE95d, Ish92]. Tolerant [Bab90b, Chexx, IEE93a, ACCB93, Bab89, Bab90a, BHMR91, EKM^{+}99, Nan91, RSS93]. tool [CFH^{+}93a, TI94b]. Tools [Ano89b]. top [CLR94, CNTS93, FKS92, MGZ93, MZDG93]. Toulouse [IEE93a]. tracing [Che93, GAR^{+}93]. Track [Shr89]. tradeoffs [NUS^{+}93, UNS^{+}94]. train [EKM^{+}99]. Transaction [GHR89, Nan91, SPB88, SGM90, Spe87, SED^{+}89]. Translation [BRG^{+}89, BKW94, Ros89]. Transparent [Gol90, RSS93, RS95]. Transputer [Ano89a, Boa90]. transputers [SR89]. treatment [IMP94]. Troy [IEE94c, SS96]. Trusted [BST95, E^{+}91, FM93, Seb91b, Seb91a, BCB88, BTM88, BL89, ES90]. Twenty [HS94b, IEE93a, MSNS91, MMH93, MS95, Shr89]. Twenty-Eighth [MS95]. Twenty-Fifth [MSNS91]. Twenty-Second [Shr89]. Twenty-Seventh [HS94b]. Twenty-Sixth [MMH93]. Twenty-Third [IEE93a]. TX [IEE93c, USE91c].

UK [Ano89b, SBC^{+}94]. uniprocessors [BRE92]. universal [CCGS92, Dri92]. Unix [Imaxx, ABG^{+}86, Ano88b, Bas91, Jal92, Mor96, SJ95, Tev87b, TBG^{+}87, TTG^{+}87, USE88, USE95a, Ano88a, Ano92d, BCC^{+}91, Dri92, Rao91, Ras91, Roy93, TS89]. USA [ACM86, ACM89, ACM91, ACM92a, ACM92b, ACM93b, ACM93a, ACM93c, Amm90, Ano88b, Ano90a, Ano90c, Ano91a, Ano91b, Ano93a, Ano93c, Ano93d, Ano94, Boa90, Ciz94, GHR89, HS94b, IEE88b, IEE88c, IEE89a, IEE89b, IEE89c, IEE89d, IEE90a, IEE90b, IEE91, IEE92a, IEE92b, IEE93b, IEE93c, IEE93e, IEE93d, IEE94f, IEE94c, IEE94d, IEE94a, IEE94e, IEE94b, IEE95a, IEE95b, IEE95c, Mai93, MMR91, MSNS91, MMH93, MS95, Shr89, SS96, USE88, USE91a, USE91c, USE92a, USE92b, USE92c, USE93a, USE93b, USE93c, USE93d, USE94, USE95b]. usage [MST93]. USENIX [USE90, Ano88b, Ano90c, Ano91a, Ano93c, Ano94, USE92c]. User [Ano89a, GSR93, OT94, SP91a, SP91b, Dea93, Mac91, Mac92, McD89, Moy93, OT95, PRK95, ST93, TNML93]. User-Level [GSR93, SP91b, OT94, SP91a, Dea93, OT95]. Users [Boa90]. Using [Dea93, DBRD91, GBB93a, GBB93b, BHMR91, BM95, FKS92, KF93, Nic91, WKF^{+}92]. UT [USE95a]. Utrecht [Ano92a]. UX [RS95].

V [ACM92b]. validation [AKST93]. VAPP [Bur90]. Variants [Mor96]. Vector [Bur90]. Vehicles [Ano90a]. Venezuela [Gir94]. Vermont [USE90]. version [OKID92]. versus [PM18]. Victoria [Ano89a]. Video [SBC^{+}94]. view [BJ94]. Virtual [Bit90, BCCR91, BCF^{+}91, CRRS93, IKWS92, Joh91, Red92, Saa92, Tev87a]. Virtually [CMS90, IKWS92, WB92]. VMS [WKF^{+}92, Wie92]. vnode [LBLM90]. Vol [HS94b]. Vol.II [Shr89]. volumes [Koo93].

WA [IEE94a, USE92a]. Wailea [HS94b, MMH93, MS95]. WARASA [JM92]. Washington [Ano90a, Ano90c]. well [BCF^{+}93]. West [Ver90]. WG10.3 [Gir94]. Whitewater [Ano93d]. Who [Ano92d]. WI [Ano93d]. Window [ES90, E^{+}91]. windows [Ano92d, GMR93]. Winter [Ano90c, USE91c, USE94]. within [BJ95]. without [Lie92]. Work [HCF^{+}94]. Working [Gir94]. Workshop [ACM93c, Ano88b, Ano92b, Bd92b, CJ92, GHR89, IEE89d, IEE90b, IEE91, IEE93b, IEE94a, IEE94e, IEE95d, SBC^{+}94, USE88, USE92a, USE90]. Workstation [Ano92b, IEE89d, IEE93b, BM92, Sha91, SCSK93]. Workstations [CP97, AG95, Mal91]. World [Ciz94]. WPI [FKL91, KF93]. wrappers [KD95]. writer [Ben92]. Writing [GJ92, Dan93]. WRTP [Bd92b]. WWOS [IEE89d]. WWOS-II [IEE89d].

X [ES90, E^{+}91]. X11

[And90, GBB93a, GBB93b].

York [Amm90].

Zurich [Bur90].

References

Accetta:1986:MNK

[ABG^{+}86] Mike Accetta, Robert Baron, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. MACH: a new kernel foundation for UNIX development. Technical report, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1986. 18 pp.

URL https://www.cs.ubc.
ca/~norm/508/2009W1/mach_
usenix86.pdf.

Arevalo:1993:FSM

[ACCB93] S. Arevalo, J. Carretero, J. L.

Castellanos, and F. Barco. A fault-tolerant server on MACH. In Anonymous [Ano93b], pages 793–800. CODEN MMICDT. ISSN 0165-6074.

ACM:1986:OOO

[ACM86] ACM, editor. OOPSLA ’86.

Object-Orientated Programming Systems, Languages and Applications. Conference Proceedings. Portland, OR, USA, 29 September–2 October 1986, volume 21(11) of ACM SIG-PLAN Notices. ACM Press, New York, NY 10036, USA, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1989:APT

[ACM89] ACM, editor. ASPLOS-III Proceedings. Third International Conference on Architectural Support for Programming Languages and Operating Systems, Boston, MA, USA, April 3– 6, 1989. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-300-0. LCCN QA76.9.A73I565 1989. ACM order number 556890. Also published as Computer architecture news, v. 17, no. 2 (Apr. 1989), Operating systems review, v. 23, special issue (Apr. 1989), and SIGPLAN notices, v. 24, special issue (May 1989).

ACM:1991:TAS

[ACM91] ACM, editor. Thirteenth ACM Symposium on Operating Systems Principles, Pacific Grove, CA, USA, October 13–16, 1991, volume 25(5) of Operating Systems Review. ACM Press, New York, NY 10036, USA, 1991. CODEN OSRED8. ISSN 0163-5980. ACM order number 534910.

ACM:1992:ASC

[ACM92a] ACM, editor. ACM SIGPLAN ’92 Conference on Programming Language Design and Implementation, San Francisco, CA, USA, June 17–19, 1992, volume 27(7) of ACM SIG-PLAN Notices. ACM Press, New York, NY 10036, USA, July 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1992:FIC

[ACM92b] ACM, editor. Fifth International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS-V), Boston, MA, USA, October 12– 15, 1992, volume 27(9) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, September 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1993:ASO

[ACM93a] ACM, editor. 14th ACM Symposium on Operating Systems Principles, Ashville, NC, USA, December 5–8, 1993, volume 27(5) of Operating Systems Review. ACM Press, New York, NY 10036, USA, December 1993. CODEN OSRED8. ISSN 0163-5980.

ACM:1993:ASC

[ACM93b] ACM, editor. 1993 ACM Sigmetrics Conference on Measurement and Modeling of Computer Systems, Santa Clara, CA, USA, May 17–21, 1993, volume 21(1) of Performance Evaluation Review. ACM Press, New York, NY 10036, USA, June 1993. CO-DEN PEREDN. ISSN 0163-5999 (print), 1557-9484 (electronic).

ACM:1993:WLC

[ACM93c]  ACM, editor.  Workshop

on Languages, Compilers and Run-Time Environments for Distributed Memory Multiprocessors, Boulder, CO, USA, 30 September – 2 October 1992, volume 28(1). ACM Press, New York, NY 10036, USA, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Apiki:1995:FNW

[AG95] S. Apiki and R. Grehan. Fastest NT workstations (midrange workstations review). Byte Magazine, 20(3):115–120, 122, March 1995. CODEN BYT-EDJ. ISSN 0360-5280.

Arakawa:1993:MVR

[AKST93] H. Arakawa, D. I. Katcher, J. K. Strosnider, and H. Tokuda. Modeling and validation of the Real-Time Mach Scheduler. In ACM [ACM93b], pages 195– 206. CODEN PEREDN. ISSN 0163-5999 (print), 1557-9484 (electronic).

Ammar:1990:PII

[Amm90] R. A. Ammar, editor. Proceedings of the ISMM International Conference Parallel and Distributed Computing, and Systems, New York, NY, USA, October 10–12, 1990. Acta Press, Anaheim, CA, USA, 1990. ISBN 0-88986-162-5. LCCN QA76.9.D5 I86 1990.

Anderson:1990:IDC

[And90] David P. Anderson. Integrated digital continuous media: a framework based on Mach, X11, and TCP/IP. Report UCB/CSD 90/566, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, March 22, 1990. 22 pp.

Anonymous:1988:DOS

[Ano88a] Anonymous. Distributed operating system is expected to follow Unix. Nikkei Electronics, 456:169–178, ???? 1988. CODEN NIERE4. ISSN 0385-1680.

Anonymous:1988:UPW

[Ano88b]  Anonymous, editor.  USE-

NIX Proceedings. Workshop on UNIX and Supercomputers, Pittsburgh, PA, USA, September 26–27, 1988. USENIX, Berkeley, CA, USA, 1988.

Anonymous:1989:CAT

[Ano89a] Anonymous, editor. Centre for Advanced Technology in Telecommunications Australian Transputer and OCCAM User Group Conference Proceedings, Melbourne, Victoria, Australia, July 6–7, 1989. Centre for Adv. Technol. Telecommun, Melbourne, Vic., Australia, 1989.

Anonymous:1989:SPC

[Ano89b]  Anonymous, editor.  Soft-

ware for Parallel Computers. Exploiting Parallelism Through Software Environments, Tools, Algorithms and Application Libraries, Hammersmith, UK, June 12–15, 1989. Unicom Seminars, Uxbridge, UK, 1989.

Anonymous:1990:CTF

[Ano90a] Anonymous, editor. Computational Technology for Flight Vehicles Symposium, Washington, DC, USA, November 5–7, 1990, volume 1(2–4) of Computing Systems in Engineering. ????, ????, 1990. CODEN COSEEO. ISSN 0956-0521.

Anonymous:1990:PAE

[Ano90b] Anonymous, editor. Proceedings of the Autumn 1990 EUUG Conference, Nice, France, October 22–26, 1990. European UNIX Users Group, Buntingford, Herts, UK, 1990. ISBN 0-9513181-8-7. LCCN ????

Anonymous:1990:PWU

[Ano90c]  Anonymous, editor.  Pro-

ceedings of the Winter 1990 USENIX Conference, Washington, DC, USA, January 22– 26, 1990. USENIX, Berkeley, CA, USA, 1990.

Anonymous:1991:PSU

[Ano91a] Anonymous, editor. Proceedings of the Summer 1991 USE-NIX Conference, Nashville, TN, USA, June 10–14, 1991. USENIX, Berkeley, CA, USA, 1991.

Anonymous:1991:SIS

[Ano91b] Anonymous, editor. SEDMS II. Symposium on Experiences with Distributed and Multiprocessor Systems, Atlanta, GA, USA, March 21–22, 1991. USE-NIX, Berkeley, CA, USA, 1991.

Anonymous:1992:DCP

[Ano92a]  Anonymous, editor.  Dis-

tributed Computing, Practice and Experience Proceedings of the Autumn 1992 Open-Forum Technical Conference, Utrecht, Netherlands, Netherlands, November 25–27, 1992. EurOpen, Buntingford, Herts, UK, 1992.

Anonymous:1992:PDW

[Ano92b] Anonymous, editor. Parallel and Distributed Workstation Systems Workshop, Florence, Italy, September 1991, volume 8(1–3) of Future Generation Computer Systems. Elsevier Science Publishers, Amsterdam, The Netherlands, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:PSE

[Ano92c] Anonymous, editor. Proceedings. SHARE Europe Spring Meeting: Managing Communications in a Global Marketplace, Cannes, France, 30 March–3 April 1992. SHARE Europe (SEAS), Geneva, Switzerland, 1992.

Anonymous:1992:WEW

[Ano92d] Anonymous. Who has the edge in windows, Apple, Microsoft or Unix. Wharton Report, 171: 1–6, November 1992. CODEN WHREEK. ISSN 0950-1800.

Anonymous:1993:AIS

[Ano93a]  Anonymous, editor.  20th

Annual International Symposium on Computer Architecture ISCA ’20, San Diego, CA, USA, May 16–19, 1993, volume 21(2). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1993. CO-DEN CANED2. ISBN 0-8186-3810-9 (paper), 0-8186-3811-7 (microfiche), 0-8186-3812-5 (case). ISSN 0163-5964 (print), 1943-5851 (electronic). LCCN QA76.9.A73 I58 1993. ACM order number 415930. IEEE catalog number 93CH3284-7. IEEE Computer Society Press order number 3810-02.

Anonymous:1993:NES

[Ano93b] Anonymous, editor. Nineteenth EUROMICRO Symposium on Microprocessing and Microprogramming (EUROMI-CRO 93). Open System Design: Hardware. Software and Applications, Barcelona, Spain, September 6–9, 1993, volume 38(1–5) (or 38(2–5)??) of Microprocessing and Microprogramming. North-Holland, Amsterdam, The Netherlands, September 1993. CODEN MMICDT. ISSN 0165-6074.

Anonymous:1993:PUS

[Ano93c] Anonymous, editor. Proceedings of the USENIX Symposium on Microkernels and Other Kernel Architectures, San Diego, CA, USA, September 20–21, 1993. USENIX, Berkeley, CA, USA, 1993.

Anonymous:1993:SAM

[Ano93d]  Anonymous, editor.  Sev-

enth Annual Midwest Computer Conference. Proceedings Manual, Whitewater, WI, USA, March 26, 1993. Univ. Wisconsin-Whitewater, Whitewater, WI, USA, 1993.

Anonymous:1994:PFU

[Ano94]  Anonymous, editor.  Pro-

ceedings of the First USENIX Symposium on Operating Systems Design and Implementation (OSDI), Monterey, CA, USA, November 14–17, 1994. USENIX, Berkeley, CA, USA, 1994.

Anonymous:1995:NIS

[Ano95] Anonymous, editor. 18th National Information Systems Security Conference: October 10– 13, 1995, Baltimore Convention Center, Baltimore, Maryland, proceedings, making security real. National Institute for Standards and Technology, Gaithersburg, MD, USA, 1995. LCCN QA76.9.A25 N38 1995. Two volumes.

Bataineh:1994:ESA

[BAA94] S. Bataineh and B. Al-Asir. Efficient scheduling algorithm for divisible and indivisible tasks in loosely coupled multiprocessor systems. Software Engineering Journal, 9(1):13–18, January 1994. CODEN SEJOED. ISSN 0268-6961.

Babaoglu:1989:FTC

[Bab89] Ozalp Babaoğlu. Fault-tolerant computing based on Mach. Technical report TR 89-1032, Cornell University, Dept. of Computer Science, Ithaca, NY, USA, August 1989. 20 pp.

Babaoglu:1990:FCBa

[Bab90a] Özalp Babaoğlu. Fault-tolerant computing based on Mach. Operating Systems Review, 24(1): 27–39, January 1990. CODEN OSRED8. ISSN 0163-5980.

Babaoglu:1990:FCBb

[Bab90b] Ozalp Babaoğlu. Fault-tolerant computing based on Mach. In USENIX [USE90], pages 185– 199. LCCN QA76.9.M45 M33 1990.

Bacon:19xx:OOR

[Bacxx] David F. Bacon. OptiMach: Optimistic recovery of Mach tasks. In ????, page ?? USE-NIX, Berkeley, CA, USA, 19xx. ISBN ???? LCCN ????

Barrera:1991:FMN

[Bar91] Joseph S. Barrera, III. A fast Mach network IPC implementation. In USENIX [USE91a], pages 1–12. LCCN QA76.8.U65 U83 1991.

Basavaiah:1991:MIC

[Bas91] Muralidhar Basavaiah. Mach interprocess communication server and network server on Berkeley UNIX. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1991. viii + 91 pp.

Baumgarten:1992:SSR

[Bau92]  U. Baumgarten.  Step by

step refinement of conceptsdesign of distributed systems with OlDiLa and Mach examples. In Jammel [Jam92], pages 262–273. ISBN 3-540-55340-1 (Berlin), 0-387-55340-1 (New York). LCCN QA76.9.A73G5 1992.

Bernadat:1993:RMM

[BB93] Philippe Bernadat and David Black. Real memory Mach. In USENIX [USE93a], pages 235–251. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Banatre:1992:SFG

[BBP92] M. Banatre, Y. Belhamissi, and I. Puaut. Some features of Gothic: a distributed objectbased system. In Cabrera and Jul [CJ92], pages 117–123. ISBN 0-8186-3015-9. LCCN QA76.76.O63 I59 1992. IEEE catalog number 92TH0477-0. IEEE Computer Society Press order number 3015.

Barker:1988:ECT

[BCB88] W. C. Barker, P. Cochrane, and M. Branstad. Embedding cryptography into a Trusted Mach system. In IEEE [IEE88b], pages 379–383. ISBN 0-8186-0895-1. LCCN TL787 .A471 1988; QA76.9.A25 A39 1988. IEEE catalog number 88CH2629-5. IEEE Computer Society order number 895.

Boyer:1991:SOD

[BCC^{+}91] F. Boyer, J. Cayuela, P. Y. Chevalier, A. Freyssinet, and D. Hagimont. Supporting an object-oriented distributed system: experience with Unix, Mach and Chorus. In Anonymous [Ano91b], pages 283–299.

Bryant:1991:SSS

[BCCR91] R. Bryant, P. Carini, H.-Y. Chang, and B. Rosenburg. Supporting structured shared virtual memory under Mach. In USENIX [USE91a], pages 59– 76. LCCN QA76.8.U65 U83 1991.

Black:1991:OVM

[BCF^{+}91] D. Black, J. Carter, G. Feinberg, R. MacDonald, S. Mangalat, E. Shienbrood, J. Van Sciver, and Ping Wang. OSF/1 virtual memory improvements. In USENIX [USE91a], pages 87–103. LCCN QA76.8.U65 U83 1991.

Balter:1993:MTW

[BCF^{+}93] R. Balter, P. Y. Chevalier, A. Freyssinet, D. Hagimont, S. Lacourte, and X. Rousset de Pina. Is the microkernel technology well suited for the support of object-oriented operating systems: the Guide experience. In Anonymous [Ano93c], pages 1–11.

Bryant:1991:EDRa

[BCR91a] R. Bryant, Hung-Yang Chang, and B. Rosenburg. Experience developing the RP3 operating system. Computing Systems, 4(3):183–216, Summer 1991. CODEN CMSYE2. ISSN 0895-6340.

Bryant:1991:EDRb

[BCR91b] R. Bryant, Hung-Yang Chang, and B. Rosenburg. Experience developing the RP3 operating system. In Anonymous [Ano91b], pages 1–18.

Bryant:1991:OSS

[BCR91c] R. M. Bryant, H.-Y. Chang, and B. S. Rosenburg. Operating system support for parallel programming on RP3. IBM Journal of Research and Development, 35(5/6):617–634, September/November 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Barton-Davis:1992:ASA

[BD92a] Paul Barton-Davis. Adding scheduler activations to Mach 3.0. Technical report 92-08-03, University of Washington, Dept. of Computer Science and Engineering, Seattle, WA, USA, August 1992. 30 pp. Revised October 1992.

Boullart:1992:RPW

[Bd92b] L. Boullart and J. A. de la Puente, editors. Real-Time Programming (WRTP ’92). Preprints of the IFAC Workshop, Bruges, Belgium, June 23–26, 1992. Pergamon Press, Oxford, UK, 1992. ISBN 0-08-041894-5. LCCN QA76.54.R423 1992.

BartonDavis:1993:ASA

[BDMVL93a] P. Barton-Davis, D. McNamee, R. Vaswani, and E. D. Lazowska. Adding scheduler activations to Mach 3.0. In USE-NIX [USE93a], pages 119–136. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Barton-Davis:1993:ASA

[BDMVL93b] Paul Barton-Davis, Dylan Mc-Namee, Raj Vaswani, and Edward Lazowska. Adding scheduler activations to Mach 3.0. In USENIX [USE93a], pages 119–136. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Benson:1992:OSS

[Ben92] G. Benson. An optimal solution to the secure reader-writer problem. In IEEE [IEE92a], pages 251–258. ISBN 0-8186-2825-1. LCCN QA 76.9 A25 I34 1992. IEEE catalog number 92CH3157-5. IEEE Computer Society Press order number 2825.

Bolosky:1989:SET

[BFS89] W. J. Bolosky, R. P. Fitzgerald, and M. L. Scott. Simple but effective techniques for NUMA memory management. Operating Systems Review, 23(5):19– 31, ???? 1989. CODEN OS-RED8. ISSN 0163-5980.

Black:1989:MEH

[BGH^{+}89] David L. Black, David B. Golub, Karl Hauth, Avadis Tevanian, and Richard Sanzi. The Mach exception handling facility. ACM SIGPLAN Notices, 24(1):45–56, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:1991:MOS

[BGJ^{+}91] D. L. Black, D. B. Golub, D. P.

Julin, R. F. Rashid, R. P. Draves, R. W. Dean, A. Forin, J. Barrera, H. Tokuda, G.-R. Malan, and D. Bohman. Microkernel operating system architecture and Mach. Journal of Information Processing, 14(4): 442–453, ???? 1991. CODEN JIPRDE. ISSN 0387-6101.

Black:1992:MOS

[BGJ^{+}92] David L. Black, David B.

Golub, Daniel P. Julin, Richard F. Rashid, Richard P. Draves, Randall W. Dean, Alessandro Forin, Joseph Barrera, Hideyuki Tokuda, Gerald Malan, and David Bohman. Microkernel operating system architecture and Mach. In USENIX [USE92a], pages 11– 30. ISBN 1-880446-42-1. LCCN QAX 32.

Black:1989:CMD

[BGW89] D. L. Black, A. Gupta, and W.- D. Weber. Competitive management of distributed shared memory. In IEEE [IEE89b], pages 184–190. ISBN 0-8186-1909-0 (paperback), 0-8186-5909-2 (microfiche), 0-8186-9909-4 (case). LCCN QA75.5 .C58 1989. IEEE catalog number 89CH2686-4.

Banatre:1993:EDR

[BHM^{+}93] M. Banatre, P. Heng, G. Muller, N. Peyrouze, and B. Rochat. An experience in the design of a reliable object based system. In IEEE [IEE93d], pages 187–190. ISBN 0-8186-3330-1. LCCN QA76.58 .I54 1993. IEEE catalog number 93TH0493-7.

Banatre:1991:HDR

[BHMR91] M. Banatre, P. Heng, G. Muller, and B. Rochat. How to design reliable servers using fault tolerant micro-kernel mechanisms. In USENIX [USE91a], pages 223–231. LCCN QA76.8.U65 U83 1991.

Bhatti:1998:CSC

[BHSC98] Nina T. Bhatti, Matti A. Hiltunen, Richard D. Schlichting, and Wanda Chiu. Coyote: a system for constructing fine-grain configurable communication services. ACM Transactions on Computer Systems, 16(4):321–366, November 1998. CODEN AC-SYEC. ISSN 0734-2071.

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tocs/1998-16-4/p321-bhatti/

.

Bitar:1990:MVM

[Bit90] Nawaf Bitar. Mach Virtual Memory Internals. UNIX-AT&T, Washington, DC, USA, 1990. various pp.

Bitar:1992:MOC

[Bit92] Nawaf Bitar. Mach: the OSF/1 core kernel: November 2-4, 1992, Engineering 819.238: lecture notes. University of California, Los Angeles, University Extension Dept. of Business, Engineering, and Management, Short Course Program, Los Angeles, CA, USA, 1992. various pp.

Bodorik:1994:MAC

[BJ94] P. Bodorik and D. N. Jutla.

Multi-view access control memory computer system. In Cizmar [Ciz94], pages 241–248. ISBN 0-89791-634-4. LCCN QA76.7 A849 1994.

Bollella:1995:SRC

[BJ95] G. Bollella and K. Jeffay. Support for real-time computing within general purpose operating systems-supporting coresident operating systems. In IEEE [IEE95c], pages 4–14. ISBN 0-8186-6980-2. LCCN QA76.54.S95 1995. IEEE catalog number 95TH8055.

Boykin:1993:PUM

[BKLL93] Joseph Boykin, David Kirschan, Alan Langerman, and Susan LoVerso. Programming under Mach. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52739-1. xvii + 490 pp. LCCN QA76.8.N49 P76 1993.

Bala:1994:SPC

[BKW94] K. Bala, M. Frans Kaashoek, and W. E. Weihl. Software prefetching and caching for translation lookaside buffers. In Anonymous [Ano94], pages 243–253.

Branstad:1989:ATM

[BL89] M. Branstad and J. Landauer. Assurance for the Trusted Mach operating system. In IEEE [IEE89a], pages 103– 108. IEEE catalog number 89CH2656-7.

Board:1990:PPN

[BL90a] J. A. Board, Jr. and J. Shue-Jen Lu. Performance of parallel neural network simulations. In Board [Boa90], pages 185–200.

Boykin:1990:MCA

[BL90b] J. Boykin and A. Langerman. Mach/4.3BSD: a conservative approach to parallelization. Computing Systems, 3(1): 69–99, Winter 1990. CODEN CMSYE2. ISSN 0895-6340.

Boykin:1990:MAC

[BL90c] Joseph Boykin and Alan Langerman. Mach/4.3BSD: a conservative approach to parallelization. Computing Systems, 3(1):69–100, 1990. CO-DEN CMSYE2. ISSN 0895-6340.

Black:1988:MEH

[Bla88] David L. Black. The Mach exception handling facility. Research paper CMU-CS-88-129, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1988. 12 pp.

Black:1990:SSCa

[Bla90a] David L. Black. Scheduling support for concurrency and parallelism in the Mach operating system. Research paper CMU-CS-90-125, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, April 1990. 16 pp. A revised version of this paper appears in the May 1990 issue of IEEE Computer [Bla90b].

Black:1990:SSCb

[Bla90b] David L. Black. Scheduling support for concurrency and parallelism in the Mach operating system. Computer, 23 (5):35–43, May 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [Bla90a].

Black:1991:PPP

[Bla91] D. L. Black. Processors, priority, and policy: Mach scheduling for new environments. In USENIX [USE91c], pages 1–12.

Bolinger:1991:PSH

[BM91] D. Bolinger and S. Mangalat. Parallelizing signal handling and process management in OSF/1. In USE-NIX [USE91a], pages 105–122. LCCN QA76.8.U65 U83 1991.

Bisiani:1992:DHW

[BM92] R. Bisiani and O. Martin.

A distributed-memory, highperformance workstation. In Anonymous [Ano92b], pages 83–91. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bryce:1995:MMM

[BM95] G. Bryce and G. Muller. Matching micro-kernels to modern applications using finegrained memory protection. In IEEE [IEE95d], pages 272–279. ISBN 0-8186-7106-8. LCCN QA76.54.I59 1995. IEEE catalog number 95TB100002.

Benson:1996:DMS

[BO96] G. D. Benson and R. A. Olsson. The design of microkernel support for the SR concurrent programming language. In Szymanski and Sinharoy [SS96], pages 227–240. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Board:1990:TRA

[Boa90] J. A. Board, editor. Transputer Research and Applications 2. NATUG-2 Proceedings of the North American Transputer Users Group, Durham, NC, USA, October 18–19, 1989. IOS Press, Amsterdam, The Netherlands, 1990.

Bolmarcich:1989:IME

[Bol89] A. S. Bolmarcich. An introduction to Mach/EPEX. Research report RC 14369 (#64355), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1989. 13 pp.

Bershad:1992:FME

[BRE92] B. N. Bershad, D. D. Redell, and J. R. Ellis. Fast mutual exclusion for uniprocessors. In ACM [ACM92b], pages 223–233. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:1989:TLB

[BRG^{+}89] D. L. Black, R. F. Rashid, D. B.

Golub, C. R. Hill, and R. V. Baron. Translation lookaside buffer consistency: a software approach. In ACM [ACM89], pages 113–122. ISBN 0-89791-300-0. LCCN QA76.9.A73I565 1989. ACM order number 556890. Also published as Computer architecture news, v. 17, no. 2 (Apr. 1989), Operating systems review, v. 23, special issue (Apr. 1989), and SIG-PLAN notices, v. 24, special issue (May 1989).

Brown:1997:LMM

[Bro97] Victoria L. Brown. Linux?

on the Macintosh? with Mach? Linux Journal, 37:??, May 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/2093.html.

Baron:1985:MOE

[BRS^{+}85] Robert Baron, Richard Rashid, Ellen Siegel, Avadis Tevanian, and Michael Young. Mach-1: An operating environment for large-scale multiprocessor applications. IEEE Software, 2 (4):65–67, July 1985. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bryant:1988:RPC

[Bry88] R. Bryant. The RP3 parallel computing environment. In Anonymous [Ano88b], pages 69–92.

Bitar:1991:MAI

[BS91] Nawaf Bitar and Eric R. Shienbrood, editors. Mach, architecture and implementation: November 4-7, 1991, Engineering 819.221: lecture notes. University of California, Los Angeles, University Extension Dept. of Business, Engineering, and Management, Short Course Program, Los Angeles, CA, USA, 1991.

Benzel:1995:ISO

[BST95] T. C. V. Benzel, E. J. Sebes, and H. Tajalli. Identification of subjects and objects in a trusted extensible client server architecture. In Anonymous [Ano95], pages 83–99. LCCN QA76.9.A25 N38 1995. Two volumes.

Bevier:1992:OSC

[BT92] W. R. Bevier and T. Taylor. Operating system correctness is a mission critical requirement. In A. K. Agrawala, K. D. Gordon, and P. Hwang, editors, Mission critical operating systems, pages 21–23. IOS Press, Amsterdam, The Netherlands, 1992.

Baron:1988:MKI

[BTG^{+}88] Robert V. Baron, Avadis Tevanian, David Golub, Richard Rashid, William Bolosky, Richard P. Draves, Jonathan Chew, David Black, and Michael Young. MACH kernel interface manual. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1988. 81 pp.

Branstad:1988:SIT

[BTM88] M. Branstad, H. Tajalli, and F. Mayer. Security issues of the Trusted Mach system. In IEEE [IEE88b], pages 362–367. ISBN 0-8186-0895-1. LCCN TL787 .A471 1988; QA76.9.A25 A39 1988. IEEE catalog number 88CH2629-5. IEEE Computer Society order number 895.

Branstad:1989:AMM

[BTMD89] M. Branstad, H. Tajalli, F. Mayer, and D. Dalva. Access meditation in a message passing kernel. In IEEE [IEE89c], pages 66–72. ISBN 0-8186-1939-2. LCCN QA 76.9 A25 I43 1989. IEEE catalog number 89CH2703-7.

Burkhart:1990:CIJ

[Bur90] H. Burkhart, editor. CON-PAR 90-VAPP IV. Joint International Conference on Vector and Parallel Processing. Proceedings, Zurich, Switzerland, September 10–13, 1990.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-53065-7. LCCN QA76.58.J65 1990.

CaraDonna:1993:LPA

[Car93] Joseph P. CaraDonna. A lock performance analysis of the Mach 3.0 micro-kernel. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1993. 156 pp.

Caswell:1989:IMD

[CB89] Deborah L. Caswell and David L. Black. Implementing a Mach debugger for multithreaded applications. Research paper CMU-CS-89-154, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, November 1989. 13 pp. To appear in the Conference Proceedings of Winter 1990 USENIX Technical Conference and Exhibition, Washington, DC, January, 1990.

Caswell:1990:IMD

[CB90] D. Caswell and D. Black. Implementing a Mach debugger for multithreaded applications. In Anonymous [Ano90c], pages 25–39.

Chen:1993:IOS

[CB93] J. B. Chen and B. N. Bershad. The impact of operating system structure on memory system performance. In ACM [ACM93a], pages 120–133. CO-DEN OSRED8. ISSN 0163-5980.

Cohn:1992:UDP

[CCGS92] D. L. Cohn, M. R. Casey, P. M.

Greenawalt, and J. E. Saldanha. A universal distributed programming paradigm for multiple operating systems. In USENIX [USE92b], pages 191– 203.

Carter:1993:FTB

[CFH^{+}93a] J. B. Carter, B. Ford, M. Hibler, R. Kuramkote, J. Law, J. Lepreau, D. B. Orr, L. Stoller, and M. Swanson. FLEX: a tool for building efficient and flexible systems. In IEEE [IEE93b], pages 198–202. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

Chevalier:1993:ESO

[CFH^{+}93b] P. Y. Chevalier, A. Freyssinet, D. Hagimont, S. Krakowiak, S. Lacourte, and X. Rousset de Pina. Experience with shared object support in the Guide system. In Anonymous [Ano93c], pages 157–173.

Canetti:1991:PCP

[CFK^{+}91] R. Canetti, L. P. Fertig, S. A.

Kravitz, D. Malki, R. Y. Pinter, S. Porat, and A. Teperman. The parallel C (pC) programming language. IBM Journal of Research and Development, 35(5/6):727–741, September/ November 1991. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Chandranmenon:1994:PEE

[Cha94] Girish P. Chandranmenon. Providing an execution environment for C* programs on a Mach-based PC cluster. Thesis (m.s.), University of New Hampshire, Durham, NH 03824, USA, 1994. ix + 34 pp.

Chen:1993:SMS

[Che93] J. B. Chen. Software methods for system address tracing. In IEEE [IEE93b], pages 178–185. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

Chen:19xx:BFS

[Chexx] Rong Chen. Building A faulttolerant system based on Mach. In ????, page ?? USE-NIX, Berkeley, CA, USA, 19xx. ISBN ???? LCCN ????

Cizmar:1994:AAC

[Ciz94] Dawn Cizmar, editor. 22nd Annual 1994 ACM Computer Science Conference. Scaling Up: Meeting the Challenge of Complexity in Real-World Computing Applications. Proceedings, Phoenix, AZ, USA, March 8– 10, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-634-4. LCCN QA76.7 A849 1994.

Ching:1991:EAP

[CJ91] W.-M. Ching and D. Ju. Execution of automatically parallelized APL programs on RP3. IBM Journal of Research and Development, 35(5/6):

767–777, September/November 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cabrera:1992:PSI

[CJ92] Luis-Felipe Cabrera and Eric Jul, editors. Proceedings of the Second International Workshop on Object Orientation in Operating systems, Dourdan, France, September 24–25, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3015-9. LCCN QA76.76.O63 I59 1992. IEEE catalog number 92TH0477-0. IEEE Computer Society Press order number 3015.

Ju:1991:EAD

[cJmC91] Dz ching Ju and Wai mee Ching. Exploitation of APL data parallelism on a sharedmemory MIMD machine. In ACM [ACM91], pages 61–72. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ACM order number 534910.

Chen:1993:PMS

[CJMT93] P.-F. Chen, J.-Y. Juang, G.-K.

Ma, and R.-L. Tsai. A prioritybased multi-class service TCP/ IP on Mach microkernel. Asia-Pacific Engineering Journal, Part A [Electrical Engineering], 3(3–4):347–363, September– December 1993. CODEN APE-JEM. ISSN 0129-5411.

Carr:1993:DC

[CKS93] H. Carr, R. Kessler, and M. Swanson. Distributed C++. In ACM [ACM93c], page 81. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chase:1994:SPS

[CLFL94] J. S. Chase, H. M. Levy, M. J. Feeley, and E. D. Lazowska. Sharing and protection in a single-address-space operating system. ACM Transactions on Computer Systems, 12(4):271– 307, November 1994. CODEN ACSYEC. ISSN 0734-2071.

Clamen:1990:RDC

[CLNW90] S. M. Clamen, L. D. Leibengood, S. M. Nettles, and J. M. Wing. Reliable distributed computing with Avalon/Common Lisp. In IEEE [IEE90a], pages 169–179. ISBN 0-8186-2036-6. LCCN QA76.7 .I576 1990. IEEE catalog number 90CH2854-8. IEEE Computer Society Press order number 2036.

Card:1994:DMD

[CLR94] R. Card, H. Le Van Gong, and P.-G. Raverdy. Design of the Masix distributed operating system on top of the Mach micro-kernel. In Girault [Gir94], pages 277–286. CODEN ITATEC. ISBN 0-444-81870-7. ISSN 0926-5473. LCCN QA76.58 .I4447 1994.

Chao:1990:MVA

[CMS90] Chia Chao, Milon Mackey, and Bart Sears. Mach on a virtually addressed cache architecture. In USENIX [USE90], pages 31– ?? LCCN QA76.9.M45 M33 1990.

Chen:1992:MSC

[CN92] Rong Chen and T. P. Ng. Microkernel support for checkpointing. In Anonymous [Ano92a], pages 35–43.

Castro:1993:MDO

[CNTS93] M. Castro, N. Neves, P. Trancoso, and P. Sousa. MIKE: a distributed object-oriented programming platform on top of the Mach microkernel. In USE-NIX [USE93c], pages 253–272. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Cabillic:1997:SEP

[CP97] Gilbert Cabillic and Isabelle Puaut. Stardust: An environment for parallel programming on networks of heterogeneous workstations. Journal of Parallel and Distributed Computing, 40(1):65–80, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/
10.1006/jpdc.1996.1271/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1271/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1271/production/ref.

CaraDonna:1993:MLP

[CPW93] J. P. CaraDonna, N. Paciorek, and C. E. Wills. Measuring lock performance in multiprocessor operating system kernels. In Anonymous [Ano93c], pages 37–56.

Chang:1992:EPMb

[CR92a] H. H. Y. Chang and B. Rosenburg. Experience porting Mach to the RP3 large-scale sharedmemory multiprocessor. In Ishida [Ish92], pages 259–267. CODEN FGSEVI. ISBN 0-444-88937-X. ISSN 0167-739X (print), 1872-7115 (electronic). LCCN QA75.I49 1990.

Chang:1992:EPMa

[CR92b] Henry H. Y. Chang and Bryan Rosenburg. Experience porting Mach to the RP3 large-scale shared-memory multiprocessor. Future Generation Computer Systems, 7(2–3):259–267, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Crandall:1990:NCP

[Cra90] Richard E. Crandall. The NeXT computer as physics machine. Computers in Physics, 4 (2):132–141, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Chew:1993:KSR

[CRRS93] Khien-Mien Chew, A. J. Reddy, T. H. Romer, and A. Silberschatz. Kernel support for recoverable-persistent virtual memory. In USENIX [USE93c], pages 215–234. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Dean:1992:DMK

[DA92] R. W. Dean and F. Armand.

Data movement in kernelized systems. In USENIX [USE92a], pages 243–261. ISBN 1-880446-42-1. LCCN QAX 32.

Dannenberg:1993:PMM

[Dan93] Roger B. Dannenberg. Performance measurements of the multimedia testbed on Mach 3.0: experience writing realtime device drivers, servers, and applications. Research paper CMU-CS-93-205, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, July 21, 1993. 22 pp.

Dannenberg:1994:PMM

[Dan94] Roger B. Dannenberg. Performance measurements of the Multimedia Testbed on Real-Time Mach. Research paper CMU-CS-94-141, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, April 1994. 18 pp.

Draves:1991:UCI

[DBRD91] R. P. Draves, B. N. Bershad, R. F. Rashid, and R. W. Dean. Using continuations to implement thread management and communication in operating systems. In ACM [ACM91], pages 122–136. CODEN OS-RED8. ISSN 0163-5980. ACM order number 534910.

Dean:1993:UCB

[Dea93] R. W. Dean. Using continuations to build a user-level threads library. In USENIX [USE93c], pages 137–151. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Demeure:1994:TSC

[DF94] I. Demeure and J. Farhat. Thread systems: concepts and examples (parallel programming). Technique et science informatiques : TSI, 13(6): 765–795, ???? 1994. CO-DEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Dorochevsky:1992:EPP

[DLR^{+}92] M. Dorochevsky, Liang-Liang Li, M. Reeve, K. Schuerman, and A. Veron. ElipSys: a parallel programming system based on logic. In Voronkov [Vor92], pages 469–471. CO-DEN LNCSD9. ISBN 3-540-55727-X (Berlin), 0-387-55727-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 1992.

Draves:1991:PRR

[Dra91] R. P. Draves. Page replacement and reference bit emulation in Mach. In USE-NIX [USE91a], pages 201–212. LCCN QA76.8.U65 U83 1991.

Draves:1992:M

[Dra92] Richard Draves. Mach. In USENIX [USE92a], pages 11– 30. ISBN 1-880446-42-1. LCCN QAX 32.

Dripke:1992:MUU

[Dri92] A. Dripke. Mach—a universal Unix kernel. object oriented micro-kernel as operating system basis. Elektronik, 41(8): 80, 82–84, April 1992. CODEN EKRKAR. ISSN 0013-5658.

Duchamp:1991:ETR

[Duc91] Dan Duchamp. Experience with threads and RPC in Mach. In USENIX [USE91b], pages 87–104. LCCN QA76.5 .S948 1991.

Drummond:1995:LLA

[DW95] J. Drummond and M. Wu. A low level analysis of the realtime Mach distributed operating system. In IEEE [IEE95c], pages 46–47. ISBN 0-8186-6980-2. LCCN QA76.54.S95 1995. IEEE catalog number 95TH8055.

Epstein:1991:PSA

[E^{+}91] Jeremy Epstein et al. A prototype B3 Trusted X Window System. In Proceedings of the Seventh Annual Computer Security Applications Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 1991. The architecture for TRW’s high assurance multi-level secure X prototype. See [ES90].

Echtle:1994:DCF

[EHP94] Klaus Echtle, D. Hammer, and David Powell, editors. Dependable computing — EDCC-1: first European Dependable Computing Conference, Berlin, Germany, October 4–6, 1994: proceedings, volume 852 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-58426-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 E33 1994.

Egan:1999:FTR

[EKM^{+}99] A. Egan, D. Kutz, D. Mikulin, R. Melhem, and D. Mossé. Fault-tolerant RT-Mach (FT-RT-Mach) and an application to real-time train control. Software—Practice and Experience, 29(4):379–395, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
55001841;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=55001841&
PLACEBO=IE.pdf.

Epstein:1990:PUM

[ES90] Jeremy Epstein and Marvin Shugerman. A Trusted X Window System Server for Trusted Mach. In Proceedings of the USENIX Mach Conference. ????, Burlington, VT, USA, October 1990. This paper describes the initial architecture of the Trusted X Window System prototype developed at TRW. This paper was superseded by the paper at the Seventh Annual Computer Security Applications Conference [E^{+}91].

Farncombe:1989:EM

[Far89]  M. Farncombe.  Equus—

computing on the move. In Anonymous [Ano89b], pages 95–96.

Forin:1991:ISMb

[FGB91a] Alessandro Forin, David Golub, and Brian Bershad. An I/O system for Mach 3.0. In USE-NIX [USE91a], pages 163–176. LCCN QA76.8.U65 U83 1991.

Forin:1991:ISMa

[FGB91b] Alessandro Forin, David Golub, and Brian N. Bershad. An I/O system for Mach 3.0. Research paper CMU-CS-91-191, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, October 1991. 11 pp.

Finkel:1992:DBM

[FKJ^{+}92] David Finkel, Robert Kinicki, Aju John, Bradford Nichols, and Somesh S. Rao. Developing benchmarks to measure the performance of the Mach operating system. Technical report 92-1, Worcester Polytechnic Institute, Computer Science Dept., Worcester, MA, USA, January 1992. 18 pp.

Finkel:1991:OWB

[FKL91] D. Finkel, R. E. Kinicki, and J. A. Lehmann. An overview of the WPI Benchmark Suite. Performance evaluation review: a quarterly publication of the Special Interest Committee on Measurement and Evaluation, 19(2):33–35, August 1991. CO-DEN PEREDN. ISSN 0163-5999 (print), 1557-9484 (electronic).

Fujinaga:1992:IMF

[FKS92] M. Fujinaga, T. Kato, and K. Suzuki. An implementation method of IN functional entities on top of distributed operating system and its performance evaluation using experimental system. IEICE Transactions on Communications, E75-B(10):1043–1051, October 1992. CODEN ITCMEZ. ISSN 0916-8516.

Ford:1994:EMM

[FL94] Bryan Ford and Jay Lepreau. Evolving Mach 3.0 to A migrating thread model. In USENIX [USE94], pages 97–114. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

Fine:1993:ADT

[FM93] T. Fine and S. E. Minear. Assuring Distributed Trusted Mach. In IEEE [IEE93c], pages 206–217. ISBN 0-7803-1298-8. LCCN TJ 217 I11c 1993. Four volumes. IEEE catalog number 93CH3307-6.

Forin:1988:DIP

[For88]  Alessandro Forin.  Design,

implementation, and performance evaluation of a distributed shared memory server for Mach. Research paper CMU-CS-88-165, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1988. v + 22 pp.

Fukuda:1993:POS

[Fuk93] A. Fukuda. Parallel operating systems. Joho-Shori (J. Information Processing Soc. Japan), 34(9):1139–1149, September 1993. CODEN JOSHA4. ISSN 0447-8053.

Furht:1994:NGR

[Fur94] B. Furht. Next generation of real-time operating systems: industrial prospective. In Halang and Stoyenko [HS94a], pages 595–596. ISBN 3-540-57558-8. LCCN QA76.54.R4216 1994.

Gruber:1991:EEO

[GADV91] O. Gruber, L. Amsaleg, L. Daynes, and P. Valduriez. Eos, an environment for objectbased systems. In Milutinovic et al. [MSNS91], pages 757– 768 (vol. 1). ISBN 0-8186-2420-5. LCCN ???? Four volumes. IEEE catalog number 91TH0394-7.

Grimsrud:1993:BHM

[GAR^{+}93] K. Grimsrud, J. Archibald, M. Ripley, K. Flanagan, and B. Nelson. BACH: a hardware monitor for tracing microprocessor-based systems. Microprocessors and Microsystems, 17(8):443–459, October 1993. CODEN MIMID5. ISSN 0141-9331.

Ginsberg:1993:UMCa

[GBB93a] Michael Ginsberg, Robert V. Baron, and Brian N. Bershad. Using the Mach communication primitives in X11. Research paper CMU-CS-93-121, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, March 1993. 8 pp.

Ginsberg:1993:UMCb

[GBB93b] Michael Ginsberg, Robert V. Baron, and Brian N. Bershad. Using the Mach communication primitives in X11. In USENIX [USE93a], pages 103–110. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Golub:1991:MDM

[GD91] D. B. Golub and R. P. Draves. Moving the default memory manager out of the Mach kernel. In USE-NIX [USE91a], pages 177–188. LCCN QA76.8.U65 U83 1991.

Giraud:1992:DOS

[GGDD92] M. Giraud, M. Gabassi, I. Demeure, and B. Dupouy. Distributed operating systems — present and future. Technical Report 93NJ00003, EDF— Electricité de France, Clamart, France, December 1992. 21 pp.

Gawlick:1989:HPT

[GHR89] D. Gawlick, M. Haynie, and A. Reuter, editors. High Performance Transaction Systems. 2nd International Workshop Proceedings, Pacific Grove, CA, USA, September 28–30, 1987. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51085-0. LCCN ????

Girault:1994:APD

[Gir94] Claude Girault, editor. Applications in Parallel and Distributed Computing. IFIP WG10.3 Working Conference, Caracas, Venezuela, April 18– 22, 1994, volume A-44 of IFIP Transactions. A. Computer Science and Technology. North-Holland, Amsterdam, The Netherlands, 1994. CODEN ITATEC. ISBN 0-444-81870-7. ISSN 0926-5473. LCCN QA76.58 .I4447 1994.

Guedes:1991:OIM

[GJ91] P. Guedes and D. P. Julin.

Object-oriented interfaces in the Mach 3.0 multi-server system. In IEEE [IEE91], pages 114–117. ISBN 0-8186-2265-2. LCCN QA76.64 .I56 1991. IEEE catalog number 91TH0392-1. IEEE Computer Society Press order number 2265.

Guedes:1992:WCA

[GJ92] P. Guedes and D. Julin. Writing a client-server application in C++. In USENIX [USE92c], pages 279–293.

Guttman:1994:TAF

[GJ94] J. D. Guttman and D. M. Johnson. Three applications of formal methods at MITRE. In Naftalin et al. [NDB94], pages 55–65. CO-DEN LNCSD9. ISBN 0-387-58555-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 1994.

Gu:1994:RRI

[GKK94] G. Gu, B. Krogh, and J. Kindy. R-TICS: a real-time infrastructure for control systems. In IEEE [IEE94c], pages 358–363. ISBN 0-8186-6510-6. LCCN TS155.63.I57 1994.

Golub:1993:MER

[GMR93] D. B. Golub, R. Manikundalam, and F. L. Rawson. MVM—an environment for running multiple DOS, Windows and DPMI programs on the microkernel. In USENIX [USE93c], pages 173–190. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Gheith:1994:KKS

[GMSS94] A. Gheith, B. Mukherjee, D. Silva, and K. Schwan. KTK: kernel support for configurable objects and invocations. In IEEE [IEE94e], pages 92–103. ISBN 0-8186-5390-6. LCCN QA76.9.D5I595 1994. IEEE catalog number 94TH0651-0.

Goldberg:1990:TRM

[Gol90] Arthur Goldberg. Transparent recovery of Mach applications. In USENIX [USE90], pages 169–184. LCCN QA76.9.M45 M33 1990.

Golub:1993:ADD

[GSR93] D. B. Golub, G. G. Sotomayor, and F. L. Rawson. An architecture for device drivers executing as user-level tasks. In USE-NIX [USE93b], pages 153–172. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Garsden:1990:CMI

[GW90a] H. Garsden and A. L. Wendelborn. A comparison of microtasking implementations of the applicative language SISAL. In Burkhart [Bur90], pages 697– 708. ISBN 3-540-53065-7. LCCN QA76.58.J65 1990.

Gould:1990:MSD

[GW90b] E. Gould and B. White. The 2.6 MSD software development environment. In Anonymous [Ano90b], pages 65–70. ISBN 0-9513181-8-7. LCCN ????

Hagimont:1994:PGO

[Hag94] D. Hagimont. Protection in the Guide object-oriented distributed system. In Tokoro and Pareschi [TP94], pages 280–298. ISBN 3-540-58202-9. LCCN QA76.64.E95 1994.

Hagimont:1994:PSO

[HCF^{+}94] Daniel Hagimont, P.-Y. Chevalier, A. Freyssinet, S. Krakowiak, S. Lacourte, J. Mossière, and X. Rousset de Pina. Persistent shared object support in the Guide system: Evaluation and related work. ACM SIGPLAN Notices, 29(10):129–144, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoven:1991:MIS

[Hov91] Rand A. Hoven. Mach interfaces to support guest O.S. debugging. In USENIX [USE91a], pages 131–148 (or 131–147??). LCCN QA76.8.U65 U83 1991.

Halang:1994:RTC

[HS94a] Wolfgang A. Halang and Alexander D. Stoyenko, editors. Real Time Computing. Proceedings of the NATO Advanced Study Institute, Sint Maarten, Dutch Antilles, October 5–17, 1992, volume 127 of NATO ASI series. Series F, Computer and systems sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57558-8. LCCN QA76.54.R4216 1994.

Hesham:1994:PTH

[HS94b] E.-R. Hesham and B. D. Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences. Vol. II: Software Technology, Wailea, HI, USA, January 4–7, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5060-5. LCCN ???? IEEE catalog number 94TH0607-2.

IEEE:1988:DPI

[IEE88a] IEEE, editor. Digest of papers: intellectual leverage/ COMPCON Spring 88, February 29-March 4, 1988, Thirty-Third IEEE Computer Society International Conference, Cathedral Hill Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0828-5 (paperback), 0-8186-4828-7 (microfiche), 0-8186-8828-9 (hardcover). LCCN QA75.5 .C58 1988.

IEEE:1988:FAC

[IEE88b] IEEE, editor. Fourth Aerospace Computer Security Applications Conference, Orlando, FL, USA, December 12–16, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0895-1. LCCN TL787 .A471 1988; QA76.9.A25 A39 1988. IEEE catalog number 88CH2629-5. IEEE Computer Society order number 895.

IEEE:1988:PRS

[IEE88c] IEEE, editor. Proceedings.

Real-Time Systems Symposium, Huntsville, AL, USA, December 6–8, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-4894-5. LCCN QA76.54 .R43 1988. IEEE catalog number 88CH2618-7.

IEEE:1989:CPF

[IEE89a] IEEE, editor. COMPASS ’89 Proceedings of the Fourth Annual Conference on Computer Assurance. Systems Integrity, Software Safety and Process Security, Gaithersburg, MD, USA, June 19–23, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2656-7.

IEEE:1989:DPC

[IEE89b] IEEE, editor. Digest of Papers. COMPCON Spring ’89. Thirty-Fourth IEEE Computer Society International Conference: Intellectual Leverage, San Francisco, CA, USA, February 27 – March 3, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1909-0 (paperback), 0-8186-5909-2 (microfiche), 0-8186-9909-4 (case). LCCN QA75.5 .C58 1989. IEEE catalog number 89CH2686-4.

IEEE:1989:PIS

[IEE89c] IEEE, editor. Proceedings 1989 IEEE Symposium on Security and Privacy, Oakland, CA, USA, May 1–3, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1939-2. LCCN QA 76.9 A25 I43 1989. IEEE catalog number 89CH2703-7.

IEEE:1989:WOS

[IEE89d] IEEE, editor. Workstation Operating Systems: Proceedings of the Second Workshop on Workstation Operating Systems (WWOS-II), Pacific Grove, CA, USA, September 27–29, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89TH0281-6.

IEEE:1990:ICC

[IEE90a] IEEE, editor. 1990 International Conference on Computer Languages, New Orleans, LA, USA, March 12– 15, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2036-6. LCCN QA76.7 .I576 1990. IEEE catalog number 90CH2854-8. IEEE Computer Society Press order number 2036.

IEEE:1990:PIW

[IEE90b] IEEE, editor. Proceedings.

IEEE Workshop on Experimental Distributed Systems, Huntsville, AL, USA, October 11–12, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. IEEE catalog number 90TH0337-6.

IEEE:1991:PIW

[IEE91] IEEE, editor. Proceedings. 1991 International Workshop on Object Orientation in Operating Systems, Palo Alto, CA, USA, October 17–18, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2265-2. LCCN QA76.64 .I56 1991. IEEE catalog number 91TH0392-1. IEEE Computer Society Press order number 2265.

IEEE:1992:PIC

[IEE92a] IEEE, editor. Proceedings. 1992 IEEE Computer Society Symposium on Research in Security and Privacy, Oakland, CA, USA, May 4–6, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2825-1. LCCN QA 76.9 A25 I34 1992. IEEE catalog number 92CH3157-5. IEEE Computer Society Press order number 2825.

IEEE:1992:PSA

[IEE92b] IEEE, editor. Proceedings. The Sixteenth Annual International Computer Software and Applications Conference, Chicago, IL, USA, September 21–25, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3000-0. LCCN ????

IEEE:1993:DPF

[IEE93a] IEEE, editor. Digest of Papers FTCS-23 The Twenty-Third International Symposium on Fault-Tolerant Computing, Toulouse, France, June 22–24, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1993. ISBN 0-8186-3680-7. LCCN QA76.5.I58 1993. IEEE catalog number 93CH3267-2. IEEE Computer Society order number 3680-02.

IEEE:1993:PFW

[IEE93b] IEEE, editor. Proceedings.

Fourth Workshop on Workstation Operating Systems, Napa, CA, USA, October 14–15, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

IEEE:1993:PIC

[IEE93c] IEEE, editor. Proceedings of the 32nd IEEE Conference on Decision and Control, San Antonio, TX, USA, December 15– 17, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1298-8. LCCN TJ 217 I11c 1993. Four volumes. IEEE catalog number 93CH3307-6.

IEEE:1993:PSI

[IEE93d] IEEE, editor. Proceedings of the Second International Conference on Parallel and Distributed Information Systems, San Diego, CA, USA, January 20–22, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3330-1. LCCN QA76.58 .I54 1993. IEEE catalog number 93TH0493-7.

IEEE:1993:PRS

[IEE93e] IEEE, editor. Proceedings. Real-Time Systems Symposium, Raleigh Durham, NC, USA, December 1–3, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4480-X. LCCN QA76.54.R434 1993. IEEE catalog number 93CH3376-1.

IEEE:1994:PIW

[IEE94a] IEEE, editor. Proceedings 11th IEEE Workshop on Real-Time Operating Systems and Software. RTOSS ’94, Seattle, WA, USA, 18–19 May 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

IEEE:1994:PSR

[IEE94b] IEEE, editor. Proceedings. 13th Symposium on Reliable Distributed Systems, Dana Point, CA, USA, October 25–27, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6575-0. LCCN QA 76.9 D5 S94 1994. IEEE catalog number 94CH35714.

IEEE:1994:PFI

[IEE94c] IEEE, editor. Proceedings of the Fourth International Conference on Computer Integrated Manufacturing and Automation Technology, Troy, NY, USA, October 10–12, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6510-6. LCCN TS155.63.I57 1994.

IEEE:1994:PIC

[IEE94d]  IEEE, editor.  Proceedings

of the International Conference on Multimedia Computing and Systems, Boston, MA, USA, May 15–19, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5530-5. LCCN QA76.575.I623 1994. IEEE catalog number 94TH0631-2.

IEEE:1994:PSI

[IEE94e] IEEE, editor. Proceedings Second International Workshop on Configurable Distributed Systems, Pittsburgh, PA, USA, March 21–23, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5390-6. LCCN QA76.9.D5I595 1994. IEEE catalog number 94TH0651-0.

IEEE:1994:PAI

[IEE94f] IEEE, editor. Proceedings the 21st Annual International Symposium on Computer Architecture, Chicago, IL, USA, April 18–21, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5510-0. ISSN 0163-5964 (print), 1943-5851 (electronic). LCCN QA76.9.A73 I58 1994. ACM order number 415940. IEEE catalog number 94CH3397-7. IEEE Computer Society order number 5510-02.

IEEE:1995:DPC

[IEE95a] IEEE, editor. Digest of Papers. COMPCON ’95. Technologies for the Information Superhighway, San Francisco, CA, USA, March 5–9, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7029-0. LCCN QA 75.5 C58 1995. IEEE catalog number 95CH35737.

IEEE:1995:PIP

[IEE95b] IEEE, editor. Proceedings 9th International Parallel Processing Symposium, Santa Barbara, CA, USA, April 25–28, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. LCCN QA 76.58 I56 1995. IEEE catalog number 95TH8052.

IEEE:1995:PRT

[IEE95c] IEEE, editor. Proceedings.

Real-Time Technology and Applications Symposium, Chicago, IL, USA, May 15–17, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6980-2. LCCN QA76.54.S95 1995. IEEE catalog number 95TH8055.

IEEE:1995:PSI

[IEE95d]  IEEE, editor.  Proceedings

Second International Workshop on Real-Time Computing Systems and Applications, Tokyo, Japan, October 25–27, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7106-8. LCCN QA76.54.I59 1995. IEEE catalog number 95TB100002.

Inouye:1992:EVA

[IKWS92] J. Inouye, R. Konuru, J. Walpole, and B. Sears. The effects of virtually addressed caches on virtual memory design and performance. Operating Systems Review, 26(4):14–29, October 1992. CODEN OSRED8. ISSN 0163-5980.

II:19xx:SOD

[Imaxx] Imag Institute. Supporting an object-oriented distributed system: experience with unix, Mach and Chorus. Technical report, Imag Institute,

                                        19xx.  URL ftp://imag.
                                        fr/pub/GUIDE/doc/07mach-
                                        chorus.ps.Z.

Issarny:1994:ETF

[IMP94] V. Issarny, G. Muller, and I. Puaut. Efficient treatment of failures in RPC systems. In IEEE [IEE94b], pages 170–180. ISBN 0-8186-6575-0. LCCN QA 76.9 D5 S94 1994. IEEE catalog number 94CH35714.

Ishida:1992:IIT

[Ish92] H. Ishida, editor. InfoJapan’90: information technology harmonizing with society: proceedings of the InfoJapan’90 Computer Conference, October 2– 5, 1990, Tokyo, Japan, volume 7(2–3) of Future Generation Computer Systems. Information Processing Society of Japan and Elsevier Science Publishers, Tokyo, Japan and Amsterdam, The Netherlands, April 1992. CODEN FGSEVI. ISBN 0-444-88937-X. ISSN 0167-739X (print), 1872-7115 (electronic). LCCN QA75.I49 1990.

Jaffer:1993:PAM

[Jaf93] Saeed Nawaz Jaffer. Performance analysis of Mach on the IBM RISC System 6000. Thesis (m.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1993. 88 pp.

Jalan:1992:CID

[Jal92] Ajay Jalan. Comparison and implementation of dynamic load balancing facilities under the UNIX and Mach operating systems. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1992. vi + 69 pp.

Jammel:1992:ARG

[Jam92] Alfons Jammel, editor. Architektur von Rechensystemen.

12. GI-ITG-Fachtagung (Architecture of Computing Systems. 12th GI-ITG-Meeting), Kiel, Germany, March 23–25, 1992. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55340-1 (Berlin), 0-387-55340-1 (New York). LCCN QA76.9.A73G5 1992.

Julin:1991:GES

[JCS^{+}91] D. P. Julin, J. J. Chew, J. M.

Stevenson, P. Guedes, P. Neves, and P. Roy. Generalized emulation services for Mach 3.0overview, experiences and current status. In USENIX [USE91a], pages 13–26. LCCN QA76.8.U65 U83 1991.

Jeffay:1994:LMT

[Jef94] K. Jeffay. On latency management in time-shared operating systems. In IEEE [IEE94a], pages 86–90. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

Jensen:1994:ADR

[Jen94] E. D. Jensen. Asynchronous decentralized realtime computer systems. In Halang and Stoyenko [HS94a], pages 347– 371. ISBN 3-540-57558-8. LCCN QA76.54.R4216 1994.

Jiang:1992:WEC

[JM92] Y. Jiang and A. Makinouchi. WARASA: an enhanced C++ for concurrent programming on shared memory multiprocessor computers. In IEEE [IEE92b], pages 257–262. ISBN 0-8186-3000-0. LCCN ????

John:1991:PEV

[Joh91] Aju John. Performance evaluation of the virtual memory management and interprocess communication features of the Mach operating system. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1991. xiv + 139 pp.

Jones:1986:MMK

[JR86] Michael B. Jones and Richard F. Rashid. MACH and Matchmaker: Kernel and language support for object-oriented distributed systems. Research paper CMU-CS-87-150, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1986. 11 pp.

Kurtzman:1995:DGO

[KD95] S. Kurtzman and K. Dattatri. Design goals of objectoriented wrappers for the Mach microkernel. In IEEE [IEE95a], pages 367–372. ISBN 0-8186-7029-0. LCCN QA 75.5 C58 1995. IEEE catalog number 95CH35737.

Kinicki:1993:CMD

[KF93] R. E. Kinicki and D. Finkel.

Comparisons of Mach distributed performance using the WPI Benchmark Suite. In Mudge et al. [MMH93], pages 40–49 (vol. 2). ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Kiczales:1993:NCO

[KLM^{+}93] G. Kiczales, J. Lamping, C. Maeda, D. Keppel, and D. NcManee. The need for customizable operating systems. In IEEE [IEE93b], pages 165–169. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

Kuechlin:1991:IMP

[KLN91] W. Kuechlin, D. Lutz, and N. Nevin. Integer multiplication in PARSAC-2 on stock microprocessors. In Mattson et al. [MMR91], pages 206–217. CO-DEN LNCSD9. ISBN 3-540-54522-0 (Berlin), 0-387-54522-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1991.

Khalidi:1993:FEP

[KN93] Y. A. Khalidi and M. N. Nelson. A flexible external paging interface. In Anonymous [Ano93c], pages 127–140.

Kawachiya:1995:EQS

[KONT95] K. Kawachiya, M. Ogata, N. Nishio, and H. Tokuda. Evaluation of QoS-Control servers on real-time mach. Lecture Notes in Computer Science, 1018:117–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koontz:1993:PBM

[Koo93] K. W. Koontz. Port buffers: a Mach IPC optimization for handling large volumes of small messages. In USENIX [USE93c], pages 89–102. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Kitayama:1993:RIE

[KTN93] Takuro Kitayama, Hideyuki Tokuda, and Tatsuo Nakajima. RT-IPC: An IPC extension for real-time Mach. In USE-NIX [USE93d], pages 91–104. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

Kupfer:1993:SM

[Kup93] Michael D. Kupfer. Sprite on Mach. In USENIX [USE93a], pages 307–322. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Lacapra:1991:MOS

[Lac91] F. Lacapra. The MACH operating system. Elettrotecnica, 78 (6):539–543, June 1991. CO-DEN ETRTAF. ISSN 0013-6131.

Langerman:1990:HMV

[LBLM90] A. Langerman, J. Boykin, S. LoVerso, and S. Mangalat. A highly-parallelized Mach-based vnode filesystem. In Anonymous [Ano90c], pages 297–312.

Lehr:1989:MMK

[Leh89] Theodore F. Lehr. MKM:

Mach Kernel Monitor description, examples and measurements. Research paper CMU-CS-89-131, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, March 1989. iii + 20 pp.

Lo:1993:ACD

[LHC93] S. L. A. Lo, N. C. Hutchinson, and S. T. Chanson. Architectural considerations in the design of real-time kernels. In IEEE [IEE93e], pages 138–147. ISBN 0-8186-4480-X. LCCN QA76.54.R434 1993. IEEE catalog number 93CH3376-1.

Lepreau:1993:ISM

[LHFL93] J. Lepreau, M. Hibler, B. Ford, and J. Law. In-kernel servers on Mach 3.0: implementation and performance. In USENIX [USE93a], pages 39–55. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Liedtke:1992:FTM

[Lie92] J. Liedtke. Fast thread management and communication without continuations. In USENIX [USE92a], pages 213–221. ISBN 1-880446-42-1. LCCN QAX 32.

Loucks:1993:MOS

[LMR93] L. Loucks, R. Manikundalam, and F. L. Rawson III. A microkernel-based operating system for personal digital assistants. In IEEE [IEE93b], pages 9–13. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

MacLachlan:1991:CCL

[Mac91] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-91-108, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, February 1991. vi + 168 pp. This is a revised version of Technical Report CMU-CS-87-156.

MacLachlan:1992:CCL

[Mac92] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-92-161, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, July 1992. v + 142 pp. Supersedes Technical Reports CMU-CS-87-156 and CMU-CS-91-108.

Maitan:1993:ETH

[Mai93] Jacek Maitan, editor. Enabling Technologies for High-Bandwidth Applications, Boston, MA, USA, September 10–11, 1992, volume 1785 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1993. CODEN PSISDG. ISBN 0-8194-0964-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK5105.5 .E44 1993.

Mallal:1991:EMP

[Mal91] Rizwan Mallal. Emulation of the Mach PMAP module for the MC88200 MMU on the Silicon Graphics Iris workstations. Thesis (m. s.), University of Vermont, Burlington, VT, USA, 1991. vi + 40 pp.

Mayer:1988:IRB

[May88] F. L. Mayer. An interpretation of a refined Bell-La Padula model for the TMach kernel. In IEEE [IEE88b], pages 368–378. ISBN 0-8186-0895-1. LCCN TL787 .A471 1988; QA76.9.A25 A39 1988. IEEE catalog number 88CH2629-5. IEEE Computer Society order number 895.

Maeda:1993:PSD

[MB93] C. Maeda and B. N. Bershad.

Protocol service decomposition for high-performance networking. In ACM [ACM93a], pages 244–255. CODEN OSRED8. ISSN 0163-5980.

Milojicic:1995:OSS

[MBS95] D. S. Milojicic, D. L. Black, and S. Sears. Operating system support for concurrent remote task creation. In IEEE [IEE95b], pages 486–492. ISBN 0-8186-7074-6. LCCN QA 76.58 I56 1995. IEEE catalog number 95TH8052.

McDonald:1989:CCL

[McD89] David B. McDonald. CMU Common Lisp user’s manual Mach/IBM RT edition. Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1989. v + 120 pp. Revised version of technical report CMU-CS-87-156.

Milojicic:2000:PM

[MDP^{+}00] Dejan S. Miloj́ičić, Fred Douglis, Yves Paindaveine, Richard Wheeler, and Songnian Zhou. Process migration. ACM Computing Surveys, 32(3):241– 299, 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/surveys/
                                        2000-32-3/p241-miloiic/p241-
                                        miloiic.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        surveys/2000-32-3/p241-miloiic/

.

Millard:1993:RSS

[MDRK93] B. R. Millard, P. Dasgupta, S. Rao, and R. Kuramkote. Run-time support and storage management for memorymapped persistent objects. In IEEE [IEE93c], pages 508–515. ISBN 0-7803-1298-8. LCCN TJ 217 I11c 1993. Four volumes. IEEE catalog number 93CH3307-6.

Matsuoka:1991:FPC

[MFY91] S. Matsuoka, S. Furuso, and A. Yonezawa. A fast parallel conservative garbage collector for concurrent objectoriented systems. In IEEE [IEE91], pages 87–93. ISBN 0-8186-2265-2. LCCN QA76.64 .I56 1991. IEEE catalog number 91TH0392-1. IEEE Computer Society Press order number 2265.

Milojicic:1993:ELD

[MGZ93] D. S. Milojicic, P. Giese, and W. Zint. Experiences with load distribution on top of the Mach microkernel. In Anonymous [Ano93c], pages 19–36.

Muller:1994:OSR

[MHP94] G. Muller, M. Hue, and N. Peyrouze. Operating system: results of the FTM experiment. In Echtle et al. [EHP94], pages 491–508. CO-DEN LNCSD9. ISBN 0-387-58426-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F38 E33 1994.

Milojicic:1994:LDI

[Mil94] Dejan S. Milojicic. Local Distribution: Implementation for the Mach Microkernel. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1994. ISBN 3-528-05424-7. xx + 149 pp. LCCN QA76.9.D5M55 1994. With a foreword by Jürgen Nehmer.

Minnick:1993:OIN

[Min93]  M. Minnick.  An object-

oriented interface to the NeXT sound driver. In Maitan [Mai93], pages 238–239. CO-DEN PSISDG. ISBN 0-8194-0964-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK5105.5 .E44 1993.

Minear:1995:PPC

[Min95] S. E. Minear. Providing policy control over object operations in a Mach based system. In USENIX [USE95a], pages 141– 155.

Mitchell:1991:MRC

[Mit91] David W. Mitchell. Mach resource control in OSF/1. In USENIX [USE91a], pages 123– 130. LCCN QA76.8.U65 U83 1991.

Mitchell:1993:MSR

[Mit93] Dave Mitchell. Mach symposium report. ;login: the USE-NIX Association newsletter, 18 (3):3–7, May/June 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Miyoshi:1998:RTJ

[MKT98] Akihiko Miyoshi, Takuro Kitayama, and Hideyuki Tokuda. A real-time Java server for real-time Mach. Parallel and Distributed Computing Practices, 1(2):??, ????

1998. CODEN ???? ISSN 1097-2803. URL http://

www.cs.okstate.edu/~pdcp/
vols/vol01/vol01no2abs.html#
miyoshi.

Milojicic:1997:CCS

[MLB^{+}97] Dejan S. Milojicic, Alan Langerman, David L. Black, Michelle Dominijanni, Randall W. Dean, and Steven J. Sears. Concurrency: a case study in remote tasking and distributed IPC in Mach. IEEE Concurrency, 5(2):39– 49, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/
p2039.pdf;   http://www.
computer.org/concurrency/
pd1997/p2039abs.htm.

Mudge:1993:PTH

[MMH93] T. N. Mudge, V. Milutinovic, and L. Hunter, editors. Proceeding of the Twenty-Sixth Hawaii International Conference on System Sciences, Wailea, HI, USA, January 5– 8, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Mattson:1991:AAA

[MMR91] H. F. Mattson, Teo Mora, and T. R. N. Rao, editors. Applied algebra, algebraic algorithms, and error-correcting codes: 9th International Symposium, AAECC-9, New Orleans, LA, USA, October 7– 11, 1991: proceedings, volume 539 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNCSD9. ISBN 3-540-54522-0 (Berlin), 0-387-54522-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1991.

Morin:1996:MUV

[Mor96] Richard Morin. Mach-based UNIX variants. SunExpert Magazine, 7(9):30, 32, 33, September 1996. ISSN 1053-9239. Discusses the University of Utah’s Mach 4 project (see http://www.cs.utah.edu/projects/flux/mach4/html/) the GNU Hurd project (see http://www.cs.pdx.edu/ trent/gnu/hurd/) and the commercial Mach Ten implementation that runs Mach on top of MacOS (see http://www.tenon.com/).

Moy:1993:TIU

[Moy93] Evelyn Moy. A TCP/IP user library in a Mach 3.0 system. Thesis (m.s.), University of Washington, Seattle, WA, USA, 1993. vi + 44 pp.

Mercer:1995:IIR

[MR95] C. W. Mercer and R. Rajkumar. An interactive interface and RT-Mach support for monitoring and controlling resource management. In IEEE [IEE95c], pages 134–139. ISBN 0-8186-6980-2. LCCN QA76.54.S95 1995. IEEE catalog number 95TH8055.

Malan:1991:MA

[MRGB91] G. Malan, R. Rashid, D. Golub, and R. Baron. DOS as a Mach 3.0 application. In USENIX [USE91a], pages 27–40. LCCN QA76.8.U65 U83 1991.

Mercer:1994:TPR

[MRZ94] C. Mercer, R. Rajkumar, and J. Zelenka. Temporal protection in real-time operating systems. In IEEE [IEE94a], pages 79–83. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

Mudge:1995:PTH

[MS95] T. Mudge and B. D. Shriver, editors. Proceedings of the Twenty-Eighth Hawaii International Conference on System Sciences, Wailea, HI, USA, January 3–6, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6930-6. LCCN ???? Five volumes.

Milutinovic:1991:PTH

[MSNS91] V. Milutinovic, B. D. Shriver, J. F. Nunamaker, Jr., and R. H. Sprague, Jr., editors. Proceedings of the Twenty-Fifth Hawaii International Conference on System Sciences, Kauai, HI, USA, January 7–10, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2420-5. LCCN ???? Four volumes. IEEE catalog number 91TH0394-7.

Mercer:1993:PCR

[MST93] C. W. Mercer, S. Savage, and H. Tokuda. Processor capacity reserves: an abstraction for managing processor usage. In IEEE [IEE93b], pages 129–134. ISBN 0-8186-4000-6. LCCN QA76.76.O63W667 1993. IEEE catalog number 93TH0553-8.

Mercer:1994:PCR

[MST94] C. W. Mercer, S. Savage, and H. Tokuda. Processor capacity reserves: operating system support for multimedia applications. In IEEE [IEE94d], pages 90–99. ISBN 0-8186-5530-5. LCCN QA76.575.I623 1994. IEEE catalog number 94TH0631-2.

Midorikawa:1995:INC

[MUI95] Hiroko Midorikawa, Shigeru Uchiyama, and Hajime Iizuka. An implementation of nCUBE C execution environment under Mach. Conference proceedings / IEEE Pacific Rim Conference on Communications, Computers and Signal Processing, pages 55–58, May 1995. ISSN 0893-4266.

Murase:1993:DOM

[MYS^{+}93] T. Murase, S. Yoshida, T. Sakon, Y. Maeyama, P. Halstead, and K. Chiba. Development of OSF/1-MK on MIPS architecture. Sumitomo Electric Technical Review, 36:72–76, June 1993. CODEN SETRAY. ISSN 0376-1207.

Milojicic:1993:TMT

[MZDG93] Dejan Milojicic, Wolfgang Zint, Andreas Dangel, and Peter Giese. Task migration on the top of the Mach microkernel. In USENIX [USE93a], pages 273–289. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Nangia:1991:TBF

[Nan91] Ashvini Nangia. Transaction based fault-tolerant computing in Mach. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1991. v + 67 pp.

Northcutt:1990:TPC

[NCS^{+}90] J. D. Northcutt, R. K. Clark, S. E. Shipman, D. P. Maynard, E. D. Jensen, F. D. Reynolds, and B. Dasarathy. Threads: a programming construct for reliable real-time distributed computing. In Ammar [Amm90], pages 299–304. ISBN 0-88986-162-5. LCCN QA76.9.D5 I86 1990.

Naftalin:1994:FIB

[NDB94] Maurice Naftalin, Tim Denvir, and Miquel Bertran, editors. FME ’94: industrial benefit of formal methods: Second International Symposium of Formal Methods Europe, Barcelona, Spain, October 24– 28, 1994: proceedings, volume 873 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-58555-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.D47 I593 1994.

Nichols:1991:DSU

[Nic91] Bradford Nichols. Distributed services using the Mach external memory management interface. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1991. v + 94 pp.

Nilsson:1992:PPR

[Nil92] M. Nilsson. Parallel processing research at SICS. Joho-Shori (J. Information Processing Soc. Japan), 33(4):421–426, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Nakajima:1993:IMP

[NKAT93] T. Nakajima, T. Kitayama, H. Arakawa, and H. Tokuda. Integrated management of priority inversion in Real-Time Mach. In IEEE [IEE93e], pages 120–130. ISBN 0-8186-4480-X. LCCN QA76.54.R434 1993. IEEE catalog number 93CH3376-1.

Nakajima:1991:MEM

[NM91] J. Nakajima and Y. H. Matsumoto. Multimedia/realtime extensions for the Mach operating system. In Anonymous [Ano91a], pages 183–198.

Nagle:1994:OAO

[NUMS94] D. Nagle, R. Uhlig, T. Mudge, and S. Sechrest. Optimal allocation of on-chip memory for multiple-API operating systems. In IEEE [IEE94f], pages 358–369. ISBN 0-8186-5510-0. ISSN 0163-5964 (print), 1943-5851 (electronic). LCCN QA76.9.A73 I58 1994. ACM order number 415940. IEEE catalog number 94CH3397-7. IEEE Computer Society order number 5510-02.

Nagle:1993:DTS

[NUS^{+}93] D. Nagle, R. Uhlig, T. Stanley, S. Sechrest, T. Mudge, and R. Brown. Design tradeoffs for software-managed TLBs. In Anonymous [Ano93a], pages 27–38. CODEN CANED2. ISBN 0-8186-3810-9 (paper), 0-8186-3811-7 (microfiche), 0-8186-3812-5 (case). ISSN 0163-5964 (print), 1943-5851 (electronic). LCCN QA76.9.A73 I58 1993. ACM order number 415930. IEEE catalog number 93CH3284-7. IEEE Computer Society Press order number 3810-02.

Nakajima:1992:MEM

[NYM92] Jun Nakajima, Masatomo Yazaki, and Hitoshi Matsumoto. Multimedia/realtime extensions for Mach 3.0. In USENIX [USE92a], pages 161–176 (or 161–175??). ISBN 1-880446-42-1. LCCN QAX 32.

Ogata:1992:DIH

[OKID92] K. Ogata, S. Kurihara, M. Inari, and N. Doi. The design and implementation of HoME (Smalltalk version). In ACM [ACM92a], pages 44–54. CO-DEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.S53 1992. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/143095/
                                        p44-ogata/.

Orman:1993:FGI

[OMOP93] H. Orman, E. Menze III, S. O’Malley, and L. Peterson. A fast and general implementation of Mach IPC in a network. In USENIX [USE93a], pages 75–88. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Onodera:1993:GCC

[Ono93]  T. Onodera.  A genera-

tional and conservative copying collector for hybrid objectoriented languages. Software —Practice and Experience, 23 (10):1077–1093, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oikawa:1994:URT

[OT94] S. Oikawa and H. Tokuda. User-level real-time threads. In IEEE [IEE94a], pages 7–11. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

Oikawa:1995:RDU

[OT95] S. Oikawa and H. Tokuda. Reflection of developing user-level real-time thread packages. Operating Systems Review, 29(4): 63–76, October 1995. CODEN OSRED8. ISSN 0163-5980.

Padmanaban:1995:RDS

[Pad95] Radhachandran Padmanaban.

Recoverable distributed shared memory on Mach. Thesis (m.s.), University of North Carolina at Charlotte, Charlotte, NC, USA, 1995. ix + 71 pp.

Phelan:1993:OPM

[PAO93] James M. Phelan, James W.

Arendt, and Gary R. Ormsby. An OS/2 personality on Mach. In USENIX [USE93a], pages 191–201. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Patience:1993:RSC

[Pat93] Simon Patience. Redirecting system calls in Mach 3.0, an alternative to the emulator. In USENIX [USE93a], pages 57– 73. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Pu:1990:IPS

[PC90] C. Pu and S.-W. F. Chen.

Implementation of a prototype superdatabase. In IEEE [IEE90b], pages 3–7. IEEE catalog number 90TH0337-6.

Peterson:1992:MK

[Pet92] J. L. Peterson. The Mach Kernel. In Anonymous [Ano92c], pages 571–578.

Peterson:1993:LOB

[Pet93] L. L. Peterson. Life on the OS/network boundary. Operating Systems Review, 27(2):94– 98, April 1993. CODEN OS-RED8. ISSN 0163-5980.

Pham:1991:EMD

[Pha91] T. Q. Pham. Experimental migration of a distributed application to a multithreaded environment. Technical Report HPL-91-155, Hewlett–Packard Laboratories, Palo Alto, CA, USA, October 1991. 37 pp.

Paul:1992:IRM

[PHY^{+}92] C. J. Paul, L. E. Holloway, D. Yan, J. K. Strosnider, and B. H. Krogh. An intelligent reactive monitoring and scheduling system. IEEE Control Systems Magazine, 12(3):78– 86, June 1992. CODEN IS-MAD7. ISSN 0272-1708.

Paciorek:1991:DMO

[PLL91] N. Paciorek, S. Lo Verso, and A. Langerman. Debugging multiprocessor operating system kernels. In Anonymous [Ano91b], pages 185–201.

Papadimitriou:2018:MVF

[PM18] Stergios Papadimitriou and Lefteris Moussiades. Mac OS versus FreeBSD: A comparative evaluation. Computer, 51(2): 44–53, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL https://www.computer.

org/csdl/mags/co/2018/02/
mco2018020044-abs.html.

Park:1995:IUC

[PRK95] Sang Seo Park, In Ho Ra, and Sung Jo Kim. Improving user convenience on a Mach-based distributed system. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 22(3):393– 402, March 1995. CODEN HJKHDC. ISSN 0258-9125.

Ragoonaden:1992:DOS

[Rag92] K. Ragoonaden. Developments in operating system technology. BT Technology Journal, 10(4): 85–95, October 1992. CODEN BTJUEH. ISSN 0265-0193.

Rao:1991:PCI

[Rao91] Somesh S. Rao. Performance comparison of interprocess communication in Mach and Unix. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1991. vii + 113 pp.

Rashid:1987:RAM

[Ras87] Richard F. Rashid. From RIG to ACCENT to MACH: The evolution of a network operating system. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 23 pp.

Rashid:1989:COS

[Ras89] R. Rashid. A catalyst for open systems (mach kernel). Datamation, 35(10):32–33, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Rashid:1991:MDO

[Ras91] Richard Rashid. Mach distributed operating systems/ Unix/AIX, 1991. ISBN 0-8186-2504-X (notes), 0-8186-2503-1 (video). 2 videocassettes (VHS) (163 min.).

Rashid:1989:MFO

[RBF^{+}89] R. Rashid, R. Baron, A. Forin, D. Golub, M. Jones, D. Orr, and R. Sanzi. Mach: a foundation for open systems (operating systems). In IEEE [IEE89d], pages 109–113. IEEE catalog number 89TH0281-6.

Reddy:1992:IRS

[Red92] Aram Jyothy Reddy. Implementing recovery support for virtual memory databases in Mach 3.0. Thesis (m.a.), University of Texas at Austin, Austin, TX, USA, 1992. ix + 36 pp.

Reynolds:1991:KSN

[RH91] F. Reynolds and J. Heller. Kernel support for network protocol servers. In USE-NIX [USE91a], pages 149–162. LCCN QA76.8.U65 U83 1991.

Rashid:1989:MSS

[RJO^{+}89] R. Rashid, D. Julin, D. Orr, R. Sanzi, R. Baron, A. Forin, D. Golub, and M. Jones. Mach: a system software kernel. In IEEE [IEE89b], pages 176–178. ISBN 0-8186-1909-0 (paperback), 0-8186-5909-2 (microfiche), 0-8186-9909-4 (case). LCCN QA75.5 .C58 1989. IEEE catalog number 89CH2686-4.

Rashid:1991:MA

[RMGB91] Richard Rashid, Gerald Malan, David Golub, and Robert Baron. DOS as a Mach 3.0 application. In USENIX [USE91a], pages 27–40. LCCN QA76.8.U65 U83 1991.

Reynolds:1990:TPC

[RNJ^{+}90] F. D. Reynolds, J. D. Northcutt, E. D. Jensen, R. K. Clark, S. E. Shipman, B. Dasarathy, and D. P. Maynard. Threads: a programming construct for reliable real-time distributed computing. International Journal of Mini and Microcomputers, 12(3):119–127, ???? 1990. CO-DEN IJMMDE. ISSN 0702-0481.

Robertson:1990:IOS

[Rob90] L. Robertson. Introduction to operating systems. In Verkerk [Ver90], pages 309–336.

Robbins:1994:ADS

[Rob94]  James P. Robbins.  Ar-

cade distributed shared memory and the Mach external memory management interface. M.s.c.s.e., Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN 46556, USA, 1994. iv + 28 pp. Thesis directed by David L. Cohn.

Rosenburg:1989:LTL

[Ros89]  B. S. Rosenburg.  Low-

synchronization translation lookaside buffer consistency in large-scale shared-memory multiprocessor. Operating Systems Review, 23(5):137–166, ???? 1989. CODEN OSRED8. ISSN 0163-5980.

Roscoe:1994:LNS

[Ros94] T. Roscoe. Linkage in the Nemesis single address space operating system. Operating Systems Review, 28(4):48–55, October 1994. CODEN OS-RED8. ISSN 0163-5980.

Roy:1993:UFA

[Roy93] P. J. Roy. Unix file access and caching in a multicomputer environment. In USENIX [USE93a], pages 21–37. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Rockhold:1994:OSS

[RP94] R. L. Rockhold and J. L. Peterson. Operating system support for shared memory clusters. In Hesham and Shriver [HS94b], pages 86–95. ISBN 0-8186-5060-5. LCCN ???? IEEE catalog number 94TH0607-2.

Russinovich:1995:ACM

[RS95] M. Russinovich and Z. Segall. Application-transparent checkpointing in Mach 3.0/UX. In Mudge and Shriver [MS95], pages 114–123 (vol. 1). ISBN 0-8186-6930-6. LCCN ???? Five volumes.

Russinovich:1993:ATF

[RSS93] M. Russinovich, Z. Segall, and D. P. Siewiorek. Application transparent fault management in fault tolerant Mach. In IEEE [IEE93a], pages 10–19. ISBN 0-8186-3680-7. LCCN QA76.5.I58 1993. IEEE catalog number 93CH3267-2. IEEE Computer Society order number 3680-02.

Rashid:1990:MSS

[RT90] R. F. Rashid and H. Tokuda.

Mach: a system software kernel. In Anonymous [Ano90a], pages 163–169. CODEN COSEEO. ISSN 0956-0521.

Saavedra:1992:SMM

[Saa92] Cesar Alfredo Saavedra. A semi-Markov model of the Mach virtual memory page replacement algorithm. Thesis (m.s.), University of Kansas, Computer Science, Lawrence, KS, USA, 1992. 73 pp.

Shepherd:1994:NOS

[SBC^{+}94] D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors. Network and Operating System Support for Digital Audio and Video. 4th International Workshop, NOSS-DAV ’93. Proceedings, Lancaster, UK, 3–5 November 1993, volume 846 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58404-8, 0-387-58404-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N394 1993. DM58.00.

Song:1993:PPB

[SC93] Inshik Song and Yookun Cho. Page prefetching based on fault history. In USENIX [USE93a], pages 203–213. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Stodolsky:1993:FIP

[SCB93] D. Stodolsky, J. B. Chen, and B. N. Bershad. Fast interrupt priority management in operating system kernels. In Anonymous [Ano93c], pages 105–110.

Swanson:1993:DSW

[SCSK93] M. Swanson, T. Critchlow, L. Stoller, and R. Kessler. The design of the schizophrenic workstation system. In USE-NIX [USE93a], pages 291–306. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Sebes:1991:OAD

[Seb91a] E. J. Sebes. Overview of the architecture of Distributed Trusted Mach. In USE-NIX [USE91a], pages 251–262. LCCN QA76.8.U65 U83 1991.

Sebes:1991:DTM

[Seb91b] Edward John Sebes. Distributed trusted Mach architecture. In USENIX [USE91a], pages 251–262. LCCN QA76.8.U65 U83 1991.

Spector:1989:HPD

[SED^{+}89] A. Z. Spector, J. L. Eppinger, D. S. Daniels, R. Draves, J. J. Bloch, D. Duchamp, R. F. Pausch, and D. Thompson. High performance distributed transaction processing in a general purpose computing environment. In Gawlick et al. [GHR89], pages 220–242. ISBN 3-540-51085-0. LCCN ????

Salem:1990:SMT

[SGM90] K. Salem and H. Garcia-Molina. System M: a transaction processing testbed for memory resident data. IEEE Transactions on Knowledge and Data Engineering, 2(1): 161–172, March 1990. CODEN ITKEEH. ISSN 1041-4347.

Shah:1991:IDL

[Sha91] Dhruve Shah. Implementation of a dynamic load balancing facility in a distributed workstation environment under the Mach operating system. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1991. vii + 82 pp.

Shekita:1991:HPF

[She91]  Eugene J. Shekita.  High-

Performance Implementation Techniques for Next-Generation Database Systems. Thesis (ph.d.), Computer Sciences Department, University of Wisconsin–Madison, Madison, WI, USA, May 1991. viii + 142 pp. Available as Technical Report TR 1026.

Shriver:1989:PTA

[Shr89] B. D. Shriver, editor. Proceedings of the Twenty-Second Annual Hawaii International Conference on System Sciences. Vol.II: Software Track, Kailua-Kona, HI, USA, January 3–6, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1912-0. LCCN ???? IEEE catalog number 89TH0243-6.

Stevenson:1995:MUG

[SJ95] J. Mark Stevenson and Daniel P. Julin. Mach-US: UNIX on generic OS object servers. In USENIX [USE95b], pages 119– 130. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

Sours:1997:CMI

[Sou97] Susan Pawlowski Sours. Cronus/ Mach integration. Technical Report RL-TR-97-134, Rome Laboratory, Air Materiel Command, Rome, NY, USA, 1997. vii + 142 pp.

Sechrest:1991:ULP

[SP91a] Stuart Sechrest and Yoonho Park. User-level physical memory management for Mach. Technical report CSE-TR-112-91, University of Michigan, Computer Science and Engineering Division, Dept. of Electrical Engineering and Computer Science, Ann Arbor, MI, USA, 1991. 11 pp.

Sechrest:1991:UPM

[SP91b] Stuart Sechrest and Yoonho Park. User-level physical memory management for Mach. In USENIX [USE91a], pages 189– 200. LCCN QA76.8.U65 U83 1991.

Spector:1988:CFD

[SPB88] A. Z. Spector, R. F. Pausch, and G. Bruell. CAMELOT: a flexible, distributed transaction processing system. In IEEE [IEE88a], page ?? ISBN 0-8186-0828-5 (paperback), 0-8186-4828-7 (microfiche), 0-8186-8828-9 (hardcover). LCCN QA75.5 .C58 1988.

Spector:1987:CDT

[Spe87] Alfred Z. Spector. Camelot: a distributed transaction facility for Mach and the Internet — an interim report. Research paper CMU-CS-87-129, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 36 pp.

Sane:1989:OSK

[SR89] A. Sane and S. S. S. P. Rao.

An operating system kernel for transputers compatible with Mach. In Anonymous [Ano89a], page 161.

Szymanski:1996:LCR

[SS96] Boleslaw K. Szymanski and Balaram Sinharoy, editors. Languages, Compilers and Run-Time Systems for Scalable Computers, Troy, NY, USA, May 22–24, 1995. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Savage:1993:RMT

[ST93] S. Savage and H. Tokuda. Realtime Mach timers: exporting time to the user. In USENIX [USE93c], pages 111–118. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

Stankovic:1994:ROS

[Sta94] J. A. Stankovic. Real-time operating systems. In Halang and Stoyenko [HS94a], pages 65–82. ISBN 3-540-57558-8. LCCN QA76.54.R4216 1994.

Subramanian:1991:MDP

[Sub91] I. Subramanian. Managing discardable pages with an external pager. In USENIX [USE91a], pages 77–85. LCCN QA76.8.U65 U83 1991.

Schwan:1992:TDM

[SZ92] K. Schwan and Hongyi Zhou. Toward dynamic, multiprocessor real-time threads. In Boullart and de la Puente [Bd92b], pages 149–154. ISBN 0-08-041894-5. LCCN QA76.54.R423 1992.

Schwan:1991:RT

[SZG91] K. Schwan, Hongyi Zhou, and A. Gheith. Real-time threads. Operating Systems Review, 25 (4):35–46, October 1991. CO-DEN OSRED8. ISSN 0163-

Schwan:1992:MRT

[SZG92] K. Schwan, Hongyi Zhou, and A. Gheith. Multiprocessor realtime threads. Operating Systems Review, 26(1):54–65, January 1992. CODEN OSRED8. ISSN 0163-5980.

Tevanian:1987:MTU

[TBG^{+}87] Avadis Tevanian, David Black, David Golub, Richard Rashid, Eric Cooper, and Michael Young. MACH threads and the UNIX kernel: The battle for control. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 12 pp.

Tevanian:1987:AIV

[Tev87a] Avadis Tevanian. Architectureindependent virtual memory management for parallel and distributed environments: the Mach approach. Research paper CMU-CS-88-106, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. x + 126 pp.

Tevanian:1987:MBF

[Tev87b] Avadis Tevanian. MACH: a basis for future UNIX development. Research paper CMU-CS-87-139, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 5 pp.

Takano:1995:IMM

[THKS95] Y. Takano, C. Howson, K. Konishi, and T. Sugawara. Implementation of Mach microkernelbased parallel operating system “Cenju-3/DE”. NEC Technical Journal = NEC giho, 48(12): 120–??, ???? 1995. CODEN NECGEZ. ISSN 0285-4139.

Tabata:1994:ICC

[TI94a] Y. Tabata and H. Iizuka. An implementation of Concurrent C on a distributed memory multiprocessor. Technical Reports of Seikei University, 31 (2):93–98, September 1994. CODEN SDKHEO. ISSN 0919-9888.

Tanaka:1994:SMT

[TI94b] S. Tanaka and H. Iizuka. A software monitoring tool for parallel programming with threads. Technical Reports of Seikei University, 31(2):99– 102, September 1994. CODEN SDKHEO. ISSN 0919-9888.

Tokuda:1994:DQC

[TK94] H. Tokuda and T. Kitayama. Dynamic QOS control based on real-time threads. In Shepherd et al. [SBC^{+}94], pages 114– 123. CODEN LNCSD9. ISBN 3-540-58404-8, 0-387-58404-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7881.4 .N394 1993. DM58.00.

Tevanian:1991:STA

[TMJY91] A. Tevanian, Jr., T. Matteson, D. Jaffee, and B. Yamamoto. Software technology at NeXT computer. In Anonymous [Ano91a], page 313.

Tokuda:1991:ERS

[TN91] H. Tokuda and T. Nakajima.

Evaluation of real-time synchronization in real-time mach. In USENIX [USE91a], pages 213–221. LCCN QA76.8.U65 U83 1991.

Tezuka:1995:EBC

[TN95] H. Tezuka and T. Nakajima.

Experiences with building a continuous media application on Real-Time Mach. In IEEE [IEE95d], pages 88–95. ISBN 0-8186-7106-8. LCCN QA76.54.I59 1995. IEEE catalog number 95TB100002.

Thekkath:1993:INP

[TNML93] C. A. Thekkath, T. D. Nguyen, E. Moy, and E. D. Lazowska. Implementing network protocols at user level. IEEE/ ACM Transactions on Networking, 1(5):554–565, October 1993. CODEN IEANEP. ISSN 1063-6692.

Tobe:1993:RPP

[Tob93] Y. Tobe. Researches on parallel processing at CMU. Joho-Shori (J. Information Processing Soc. Japan), 34(7):909–914, July 1993. CODEN JOSHA4. ISSN 0447-8053.

Toftner:1989:ESM

[Tof89] Elizabeth Claire Toftner. An extensible software message monitor for the Mach operating system. Thesis (m.s.), California Polytechnic State University, Pomona, CA, USA, 1989. vii + 56 pp.

Tokuda:1995:OSS

[Tok95] H. Tokuda. Operating system support for continuous media applications-RT-Mach extensions. In IEEE [IEE95d], pages 256–262. ISBN 0-8186-7106-8. LCCN QA76.54.I59 1995. IEEE catalog number 95TB100002.

Tokoro:1994:OPE

[TP94] M. Tokoro and R. Pareschi, editors. Object-Oriented Programming. 8th European Conference, ECOOP ’94. Proceedings, Bologna, Italy, July 4–8, 1994. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58202-9. LCCN QA76.64.E95 1994.

Tevanian:1989:MMF

[TS89] A. Tevanian, Jr. and B. Smith. Mach: the model for future Unix (object-oriented operating system). Byte Magazine, 14 (12):411–417, November 1989. CODEN BYTEDJ. ISSN 0360-5280.

Thompson:1990:SCN

[TS90] Tom Thompson and Ben Smith. Sizing up the Cube: The NeXT Computer – advanced features, fair performance. Byte Magazine, 15(1):169–176, January 1990. CODEN BYTEDJ. ISSN 0360-5280.

Takashio:1996:DID

[TST96] K. Takashio, H. Shitomi, and M. Tokoro. Design and implementation of DROL runtime environment on real-time Mach kernel. Lecture Notes in Computer Science, 1107:257– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tevanian:1987:UIS

[TTG^{+}87] Avadis Tevanian, Jr., Mary Thompson, David Golub, Richard Rashid, William Bolosky, Michael Young, and Richard Sanzi. A UNIX interface for shared memory and memory mapped files under MACH. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 14 pp. URL http:

//cs.cmu.edu/afs/cs.cmu.
edu/project/mach/public/
doc/published/mapfiles87.
ps.

Uhlig:1992:STM

[Uhl92] Richard Uhlig. Software TLB management in OSF/1 and Mach 3.0. Technical report CSE-TR-156-93, University of Michigan, Computer Science and Engineering Division, Dept. of Electrical Engineering and Computer Science, Ann Arbor, MI, USA, December 4, 1992. 13 pp.

Uhlig:1994:DTS

[UNS^{+}94] R. Uhlig, D. Nagle, T. Stanley, T. Mudge, S. Sechrest, and R. Brown. Design tradeoffs for software-managed TLBs. ACM Transactions on Computer Systems, 12(3):175–205, August 1994. CODEN ACSYEC. ISSN 0734-2071.

USENIX:1988:PUS

[USE88]  USENIX, editor.  Proceed-

ings. UNIX Security Workshop, Portland, OR, USA, August 29–30, 1998. USENIX, Berkeley, CA, USA, 1988.

USENIX:1990:MUW

[USE90] USENIX, editor. Mach: USE-NIX workshop proceedings: October 4–5, 1990, Burlington, Vermont. USENIX, Berkeley, CA, USA, 1990. LCCN QA76.9.M45 M33 1990.

USENIX:1991:PUM

[USE91a] USENIX, editor. Proceedings of the USENIX Mach Symposium: November 20–22, 1991, Monterey, California, USA. USE-NIX, Berkeley, CA, USA, 1991. LCCN QA76.8.U65 U83 1991.

USENIX:1991:SIS

[USE91b] USENIX, editor. SEDMS II — Symposium on Experiences with Distributed and Multiprocessor Systems: March 21–22, 1991, Atlanta, GA. USENIX, Berkeley, CA, USA, March 21– 22, 1991. LCCN QA76.5 .S948 1991.

USENIX:1991:UAP

[USE91c] USENIX, editor. USENIX Association. Proceedings of the Winter 1991 USENIX Conference, Dallas, TX, USA, January 21–25, 1991. USENIX, Berkeley, CA, USA, 1991.

USENIX:1992:PUW

[USE92a] USENIX, editor. Proceedings of the USENIX Workshop on Micro-Kernels and Other Kernel Architectures: 27–28 April, 1992, Seattle, WA, USA. USE-NIX, Berkeley, CA, USA, 1992. ISBN 1-880446-42-1. LCCN QAX 32.

USENIX:1992:SIS

[USE92b]  USENIX, editor.  SEDMS

III. Symposium on Experiences with Distributed and Multiprocessor Systems, Newport Beach, CA, USA, March 26– 27, 1992. USENIX, Berkeley, CA, USA, 1992.

USENIX:1992:UCT

[USE92c] USENIX, editor. USENIX C++ Technical Conference Proceedings, Portland, OR, USA, August 10–13, 1992. USENIX, Berkeley, CA, USA, 1992.

USENIX:1993:MIS

[USE93a] USENIX, editor. Proceedings of the USENIX Mach III Symposium, April 19–21, 1993, Santa Fe, New Mexico, USA. USE-NIX, Berkeley, CA, USA, April 19–21, 1993. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85

USENIX:1993:PUMa

[USE93b] USENIX, editor. Proceedings of the USENIX Mach III Symposium: April 19–21, 1993, Santa Fe, New Mexico, USA. USE-NIX, Berkeley, CA, USA, 1993. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

USENIX:1993:PUMb

[USE93c] USENIX, editor. Proceedings of the USENIX Mobile and Location-Independent Computing Symposium, Cambridge, MA, USA, August 2–3, 1993. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

USENIX:1993:PUS

[USE93d]  USENIX, editor.  Proceed-

ings of the USENIX Symposium on Microkernels and Other Kernel Architectures: September 20–21, 1993, San Diego, California, USA. USE-NIX, Berkeley, CA, USA, 1993. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

USENIX:1994:PWU

[USE94] USENIX, editor. Proceedings of the Winter 1994 USENIX Conference: January 17–21, 1994, San Francisco, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

USENIX:1995:UUS

[USE95a] USENIX, editor. 5th USENIX UNIX Security Symposium, June 5–7, 1995. Salt Lake City, UT. USENIX, Berkeley, CA, USA, June 5–7, 1995.

USENIX:1995:PUT

[USE95b] USENIX, editor. Proceedings of the 1995 USENIX Technical Conference: January 16–20, 1995, New Orleans, Louisiana, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

Vaughan:1992:CCA

[VBD^{+}92] F. Vaughan, T. Lo Basso, A. Dearle, C. Marlin, and C. Barter. Casper: a cached architecture supporting persistence. Computing Systems, 5(3):337–363, Summer 1992. CODEN CMSYE2. ISSN 0895-6340.

Verkerk:1990:CSC

[Ver90] C. Verkerk, editor. 1989 Cern School of Computing. Proceedings (CERN 90-06), Bad Herrenalb, West Germany, August 20 – September 2, 1989. CERN, Geneva, Switzerland, 1990.

Voelcker:1989:TS

[Voe89] J. Voelcker. Technology ’89:

software. IEEE Spectrum, 26 (1):37–39, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voronkov:1992:LPA

[Vor92] A. Voronkov, editor. Logic programming and automated reasoning: international conference, LPAR ’92, St. Petersburg, Russia, July 15–20, 1992: proceedings, volume 624 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNCSD9. ISBN 3-540-55727-X (Berlin), 0-387-55727-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63 .I55 1992.

vanRenesse:1992:RMB

[vRBC^{+}92] R. van Renesse, K. Birman, R. Cooper, B. Glade, and P. Stephenson. Reliable multicast between microkernels. In USENIX [USE92a], pages 269– 283. ISBN 1-880446-42-1. LCCN QAX 32.

Wahl:1990:OSF

[Wah90] P. Wahl. The Open Software Foundation: OSF/1. Office Management, 38(4):14–18, April 1990. CODEN OF-MADG. ISSN 0343-2319.

Wheeler:1992:CMV

[WB92] B. Wheeler and B. N. Bershad. Consistency management for virtually indexed caches. In ACM [ACM92b], pages 124– 136. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welch:1991:FSB

[Wel91] B. Welch. The file system belongs in the kernel. In USE-NIX [USE91a], pages 233–250. LCCN QA76.8.U65 U83 1991.

Wendorf:1988:IET

[Wen88] J. W. Wendorf. Implementation and evaluation of a timedriven scheduling processor. In IEEE [IEE88c], pages 172–180. ISBN 0-8186-4894-5. LCCN QA76.54 .R43 1988. IEEE catalog number 88CH2618-7.

Wolfer:1993:IIS

[WGR93] J. Wolfer, T. Grace, and J. Roberge. An investigation of image synthesis software migration to the BBN TC2000 Butterfly II Supercomputer. In Anonymous [Ano93d], pages 23–32.

Wiecek:1992:VM

[Wie92] Cheryl A. Wiecek. VMS on Mach. Operating Systems Review, 26(2):15, April 1992. CO-DEN OSRED8. ISSN 0163-5980.

Williams:1988:NAD

[Wil88] T. Williams. The NeXT architecture: design for the ’90s? Computer Design, 27(22):27– 28, December 1988. CODEN CMPDAM. ISSN 0010-4566.

Wiecek:1992:MPV

[WKF^{+}92] C. A. Wiecek, C. G. Kaler, S. Fiorelli, W. C. Davenport, Jr., and R. C. Chen. A model and prototype of VMS using the Mach 3.0 kernel. In USE-NIX [USE92a], pages 187–203. ISBN 1-880446-42-1. LCCN QAX 32.

Wilson:1993:HAD

[WLT93] A. W. Wilson, Jr., R. P. LaRowe, Jr., and M. J. Teller. Hardware assist for distributed shared memory. In IEEE [IEE93c], pages 246–255. ISBN 0-7803-1298-8. LCCN TJ 217 I11c 1993. Four volumes. IEEE catalog number 93CH3307-6.

Waldspurger:1994:LSF

[WW94] C. A. Waldspurger and W. E. Weihl. Lottery scheduling: flexible proportional-share resource management. In Anonymous [Ano94], pages 1–11.

Wendorf:1989:SOS

[WWT89] J. W. Wendorf, R. G. Wendorf, and H. Tokuda. Scheduling operating system processing on small-scale multiprocessors. In Shriver [Shr89], pages 904– 913 (vol. 2). ISBN 0-8186-1912-0. LCCN ???? IEEE catalog number 89TH0243-6.

Yep:1992:DSB

[Yep92] Christy Yep. A debugging support based on breakpoints for distributed programs running under Mach. Thesis (m.comp.sc.), Concordia University, Ottawa, ON, Canada, 1992. various pp. Available from National Library of Canada = Bibliothéque nationale du Canada, University Microfilms order no. UMI00444621.

Yuhara:1994:EPD

[YMBM94] M. Yuhara, C. Maeda, B. N.

Bershad, and J. E. B. Moss. Efficient packet demultiplexing for multiple endpoints and large messages. In USENIX [USE94], pages 153–165. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

Yoshida:1991:PSB

[YT91] H. Yoshida and T. Takahara.

A production system based on distributed artificial intelligence. Academic Reports, Faculty of Engineering, Tokyo Institute of Polytechnics, 14(1): 13–21, ???? 1991. CODEN TOKIDC. ISSN 0387-6055.

Yang:1994:AP

[YTC94] Chao-Tung Yang, Shian-Shyong Tseng, and Chang-Sheng Chen. The anatomy of parafrase-2. Proceedings of the National Science Council, Republic of China, Part A [Physical Science and Engineering], 18(5): 450–462, September 1994. CO-DEN PNAEE2. ISSN 0255-6588.

Yee:1988:SSS

[YTS88] B. S. Yee, J. D. Tygar, and A. Z. Spector. StrongBox: support for self-securing programs. In USENIX [USE88], page 50.